### v2

# Data Poisoning for In-context Learning

Abstract
In the domain of large language models (LLMs), in-context learning (ICL) has been recognized for its innovative ability to adapt to new tasks, relying on examples rather than retraining or fine-tuning. This paper delves into the critical issue of ICL’s susceptibility to data poisoning attacks, an area not yet fully explored. We wonder whether ICL is vulnerable, with adversaries capable of manipulating example data to degrade model performance. To address this, we introduce ICLPoison, a specialized attacking framework conceived to exploit the learning mechanisms of ICL. Our approach uniquely employs discrete text perturbations to strategically influence the hidden states of LLMs during the ICL process. We outline three representative strategies to implement attacks under our framework, each rigorously evaluated across a variety of models and tasks. Our comprehensive tests, including trials on the sophisticated GPT-4 model, demonstrate that ICL’s performance is significantly compromised under our framework. These revelations indicate an urgent need for enhanced defense mechanisms to safeguard the integrity and reliability of LLMs in applications relying on in-context learning.
Machine Learning, ICML
1 Introduction
In recent years, In-Context Learning (ICL) (Brown et al., 2020; Min et al., 2022) has emerged as an important component of large language models (LLMs). Unlike traditional machine learning algorithms that require extensive retraining or fine-tuning to adapt new data (Hoi et al., 2021; Zhang & Yang, 2021; Zhuang et al., 2020), ICL enables LLMs to make predictions based on extra information in the prompt (e.g. a few examples related to a specific task), without changing the model parameters. For example, consider the task of predicting a person’s nationality and the prompt consists of examples, e.g. “Albert Einstein was German; Isaac Newton was English;”, followed by the query “Thomas Edison was”, an LLM such as GPT-4 will predict “American” accurately. The efficiency and flexibility of ICL have gained significant attention and revolutionized various real-world applications, ranging from reasoning (e.g. chain-of-thought (Wei et al., 2022; Wang et al., 2022)) to retrieval-augmented-generation (RAG) (Lewis et al., 2020).
The success of ICL depends critically on the examples it utilizes. Studies have shown that the ICL performance is sensitive to certain characteristics of demonstrations, e.g., the selection of examples (Wang et al., 2023) and the order of examples in demonstration (Min et al., 2022). Therefore, it naturally raises a question: Is ICL vulnerable to potential data poisoning attacks? In reality, malicious actors may manipulate demonstration data to degrade the model performance. For example, adversaries can execute data poisoning attacks by strategically altering the examples in demonstrations used in ICL. Such attacks aim to disrupt the learning process, leading to inaccurate or biased predictions.
In this paper, we aim to answer the above question by delving into data poisoning attacks in ICL and uncovering the vulnerability of ICL against these attacks. We consider the standard pipeline of ICL, where examples are randomly selected from a certain data set to tailor ICL for a downstream task. In terms of the attack, we assume that the adversary deliberately alters some examples in the data, and his goal is to ensure that the learning process is adversely affected and the overall performance decreases. This scenario is both significant and practical. For instance, in a RAG system where demonstrations are retrieved from a dataset, the attacker manipulates some data from a specific domain or a task (e.g. reviews about a specific brand) to downgrade the quality of the generated content of LLM related to that domain/task. To the best of our knowledge, we are the first to investigate the vulnerability of ICL against data poisoning attacks.
Data poisoning in ICL faces both unique challenges specific to ICL and common obstacles in traditional data poisoning. First, in contrast to traditional learning algorithms with explicit training objectives, ICL enables LLMs to learn from demonstrations without explicit optimization (Brown et al., 2020; Min et al., 2022). Since traditional poisoning strategies are designed specifically to target the training process and exploit loss functions in conventional models (Biggio et al., 2012; Steinhardt et al., 2017; Geiping et al., 2020; He et al., 2023), the discrepancy between ICL and these strategies implies that they are not directly applicable to ICL. Conducting effective data poisoning attacks for ICL demands a thorough understanding of the unique learning mechanism of ICL. Second, similar to traditional attacking methods, data poisoning for ICL also requires creating samples that are imperceptible to humans yet disruptive. These poisoned examples must seamlessly integrate with the other data to harm the learning process in a subtle way. While both traditional data poisoning and the attack in ICL require careful designs of the attack, for ICL, one extra challenge arises from the discrete vocabulary of language models, making it hard to manipulate inputs for effective disturbance (Lei et al., 2019; Xu et al., 2023).
To tackle the above challenges, we introduce a novel and versatile attacking framework, ICLPoison, to exploit the unique learning mechanism of ICL. In particular, previous research (Xie et al., 2021; Hendel et al., 2023; Liu et al., 2023b; Wang et al., 2023) has shown a strong correlation between ICL performance and the hidden states within LLMs. Our framework, ICLPoison, ingeniously distorts these hidden states through strategic text perturbations—a non-trivial adaption given the attacker’s limited ability to only alter the examples in the demonstration. We further design three strategies for instantiating and optimizing poisoning attacks under the ICLPoison framework. Comprehensive experiments across various LLMs and tasks demonstrate the effectiveness of our methods, highlighting the vulnerability of ICL. Notably, we have successfully degraded the performance of ICL in advanced models, including GPT-4 (a 10% decrease in ICL accuracy). Our study significantly advances the understanding of ICL’s vulnerabilities to data poisoning, bolstering the security and reliability of LLMs.
2 Preliminary
In this section, we introduce some background of ICL. We then present the preliminary results that the distortion in hidden states indeed affects the ICL performance, which inspires our framework. Key notations are also introduced.
2.1 In-context Learning (ICL)
[…]
2.2 Understanding of ICL via Hidden States
The emergence of ICL ability in LLMs has gained substantial attention, particularly in understanding its working mechanisms (Xie et al., 2021; Hendel et al., 2023; Von Oswald et al., 2023; Garg et al., 2022; Bai et al., 2023). Researchers have demonstrated that during ICL, LLMs can effectively extract “latent concepts” (denoted as $\theta$) from demonstrations (Xie et al., 2021; Wang et al., 2023), and the ICL predicting procedure can be decomposed as
$P(y^{query}|[S,x^{query}])$ (1)
[…]
where $P(\cdot)$ denotes the probability distribution during the generation process of LLMs. In other words, ICL learns $\theta$ from the demonstration first and then make a prediction for $x^{query}$. The latent concept $\theta$ characterizes the intricate patterns and contextual information embedded in the data regarding the task. A key in this process is the hidden states, represented as $h$, which are defined as the representations of the last token of the input prompt at different layers of the model (as indicated by various studies (Hendel et al., 2023; Liu et al., 2023b; Todd et al., 2023)). The hidden states are used in encoding the latent concepts. In particular, consider a model $f$ composed of $L$ transformer layers, where each layer produces a vector representation $h_{l}(p,f)\in\mathbb{R}^{d}$ for the last token of the input prompt $p$ and $l\in[L]$. It has been observed that the hidden state $h_{l}$ at certain intermediate layers effectively works as $\theta$ (Hendel et al., 2023). Moreover, further research (Liu et al., 2023b; Todd et al., 2023) has been conducted to distill more concise representations from the set of hidden states $\{h_{l}\}_{l=1}^{L}$, enhancing the concept formation in ICL.
2.3 Preliminary Experiments
As discussed in the above, hidden states of the input prompt encode the latent concept and are closely related to the performance of ICL. Therefore, we conduct preliminary experiments to explore whether directly perturbing these hidden states can degrade ICL, and whether ICL is more vulnerable when perturbing the hidden states in all layers.
[…]
Figure 1: Results when directly perturbing the hidden states with Gaussian noises of various standard deviations (noise strength).
3 A Novel Framework: ICLPoison
In this section, we introduce a novel framework, ICLPoison, to conduct data poisoning by distorting hidden states of LLMs during the ICL process.
3.1 Threat Model
We assume that attackers have the full access to a portion of the dataset, specifically a subset related to a particular task. However, they may lack complete knowledge of the entire prompting set. Crucially, they are unaware of the details of the ICL process, including the test data, the Large Language Models (LLMs) employed, and specific ICL configurations like the number of examples and the templates used for demonstrations. Despite these limitations, the attackers can leverage open-source LLMs to introduce poisoned data into the subset they control.
3.2 ICLPoison Design
As highlighted in Section 2.1, ICL differs from traditional learning algorithms since it lacks an explicit training objective that can be directly targeted by data poisoning attacks. To address this unique challenge, we draw inspiration from the dynamics of hidden states from Section 2.2, which indicates that the representation of the last token of the demonstration encodes the latent concept about the task. Building on this understanding, we introduce ICLPoison, a novel data poisoning framework specifically designed for the ICL process. ICLPoison strategically alters examples in demonstrations to distort the hidden states in the LLM, leveraging these internal dynamics to achieve our poisoning goals.
The details of ICLPoison are as follows. We focus on a surrogate LLM $f$ with $L$ layers and a task-specific prompt set $D^{p}_{t}=\{(x_{i,t}^{p},y_{i,t}^{p})\}_{i=1}^{N}$ from task $t$. Our objective is to diminish the ICL accuracy for the task $t$ by maximizing the distortion of hidden states in samples from $D^{p}_{t}$. We propose perturbing the input $x^{p}_{i,t}\sim\mathcal{X}_{t}$ using a transformation $\delta:\mathcal{X}_{t}\rightarrow\mathcal{X}_{t}$ while keeping the label $y^{p}_{i,t}$ unchanged. The perturbation $\delta$ must be imperceptible to humans, hence it is constrained within a set $\Delta$ of imperceptible mappings. Details of $\delta$ will be discussed in Section 3.3. For each example $(x_{i,t}^{p},y_{i,t}^{p})$ in $D^{p}_{t}$, we extract its hidden states. Since the attacker lacks the knowledge of the test data, we use a dummy query $x^{query}_{t}\sim\mathcal{X}_{t}$ as a stand-in (as mentioned in (Hendel et al., 2023)). We then concatenate $(x_{i,t}^{p},y_{i,t}^{p})$ with $x^{query}_{t}$ to create a demonstration, and denote $h_{l}(x_{i,t}^{p},f)$ as the representation of the last token in the $l^{th}$ layer of model $f$. Since our focus is on perturbing $x^{p}_{i,t}$, we omit $y^{p}_{i,t}$ in the following discussion. The representations from all $L$ layers of model $f$ regarding input $x_{i,t}^{p}$ are denoted as $H(x_{i,t}^{p},f):=\{h_{l}(x_{i,t}^{p},f)\}_{l=1}^{L}$, representing the hidden states for $x^{p}_{i,t}$ under model $f$. For the perturbed input, the hidden states are $H(\delta_{i}(x_{i,t}^{p}),f)$, with $\delta_{i}$ being the specific perturbation for $x_{i,t}^{p}$.
Our observations in Figure 1 indicate that altering the hidden state of just one layer is not optimal to significantly impact ICL performance. Therefore, we aim to maximize the minimum difference across all layers between the original and perturbed hidden states. To normalize differences across layers with varying scales, we use the normalized $L_{2}$ norm to measure the distance of the hidden state between the original example and the perturbed one for each layer: $l_{d}(h_{l}(x_{i,t}^{p},f),h_{l}(\delta_{i}(x_{i,t}^{p}),f))=\|\frac{h_{l}(x_{i,t}^{p},f)}{\|h_{l}(x_{i,t}^{p},f)\|_{2}}-\frac{h_{l}(\delta_{i}(x_{i,t}^{p}),f)}{\|h_{l}(\delta_{i}(x_{i,t}^{p}),f)\|_{2}}\|_{2}$. The distortion between $x_{i,t}^{p}$ and $\delta_{i}(x_{i,t}^{p})$ is further defined as:
$\mathcal{L}_{d}(H(x_{i,t}^{p},f),H(\delta_{i}(x_{i,t}^{p}),f))$ (2)
$=\min\limits_{l\in[L]}l_{d}(h_{l}(x_{i,t}^{p},f),h_{l}(\delta_{i}(x_{i,t}^{p}),f)).$
The attacking objective of ICLPoison then becomes
$\max\limits_{\delta_{i}\in\Delta}\mathcal{L}_{d}(H(x_{i,t}^{p},f),H(\delta_{i}(x_{i,t}^{p}),f)).$ (3)
In other words, $\mathcal{L}_{d}(H(x_{i,t}^{p},f),H(\delta_{i}(x_{i,t}^{p}),f))$ denotes the minimum changes (or lower bound of the distortion) caused by the perturbation $\delta_{i}$ across all the layers of the model. This approach ensures that the perturbation $\delta_{i}$ introduces the most substantial change to the hidden states in the LLM during the ICL process, potentially enhancing the effectiveness of the poisoning attack.
3.3 Attacking Algorithms
The perturbation $\delta$ is a crucial component of the ICLPoison framework, designed to be imperceptible to humans while effective in manipulating the performance. These are also common requirements for NLP attacks (Ebrahimi et al., 2017; Jin et al., 2020; Xu et al., 2023; Li et al., 2018). On the other hand, the discrete nature of the objective in Eq.3 poses additional challenges in the optimization. To address these challenges as well as showcase the versatility of our framework, we introduce three representative perturbations: synonym replacement, character replacement, and adversarial suffix. These methods demonstrate the adaptability of ICLPoison across different levels of text manipulation: Synonym replacement evaluates the word-level vulnerability of ICL and subtly changes the semantics; character replacement involves minimal but precise alterations, such as changing individual letters, making it less noticeable to human reviewers (see examples in Appendix A.3); and adversarial suffix test token-level vulnerabilities in ICL, accommodating different models by adapting to their unique tokenization methods. In general, these perturbations allow for a comprehensive evaluation of the vulnerability in ICL. The optimization of these perturbations is efficiently managed through a greedy search method, proven effective in similar contexts (Lei et al., 2019; Bao et al., 2022).
Synonym Replacement. This approach involves substituting words in a text with their synonyms, aiming at preserving the semantic meaning and grammatical structure (Jin et al., 2020; Xu et al., 2023). Therefore, it serves as an appropriate example of $\delta$ in Section 3.2. Within our framework, we limit the number of word replacements (denoted as $k$) to maintain the perturbation’s imperceptibility. For a text composed of a sequence of $n$ words $x=[w_{1},...,w_{n}]$, $\delta(x)$ is defined as $[s_{1},...,s_{n}]$, where $s_{i}$ is either a synonym of $w_{i}$ (if selected for replacement) or remains as $w_{i}$ (if not replaced). Our objective is to identify the optimal replacements that maximize the objective in Eq.3. This brings natural challenges: which words are to be replaced and what synonyms are suitable. Following a strategy similar to (Jin et al., 2020), we adopt a two-step optimization process. Firstly, we calculate an importance score for each word, selecting the top-$k$ words with the highest scores for replacement. The importance score for word $w_{i}$ is computed as the distortion (in Eq.2) before and after deleting $w_{i}$, expressed as:
$I_{w_{i}}=\mathcal{L}_{d}(H(x,f),H(x_{\text{\textbackslash}w_{i}},f))$ (4)
where $x_{\text{\textbackslash}w_{i}}$ denotes the text after the removal of $w_{i}$. In the second step, we use a greedy search method. This method involves iteratively finding the best replacement for each selected word, one at a time, while keeping the other words fixed (Yang et al., 2020; Lei et al., 2019). In each iteration, we focus on one of the selected words, replacing it with its most appropriate synonym. We retrieve synonyms using GloVe word embeddings (Pennington et al., 2014), selecting those with embeddings most similar (based on cosine similarity) to the original word. After identifying potential synonyms, each synonym is temporarily substituted into the text, and the loss function in Eq.3 is evaluated. The synonym that results in the highest loss is then chosen as the final replacement. We then proceed to the next word, repeating this replacement process until traversing all selected words. We present the detailed algorithm in Algorithm. 1.
Character Replacement. This method is similar with the synonym replacement approach, but focuses on replacing individual characters instead of whole words (Ebrahimi et al., 2017; Lei et al., 2019; Xu et al., 2023). When changing only a few letters, this method can be less detectable to humans and maintain the word’s pronunciation and basic structures (Ebrahimi et al., 2017). We limit the number of character replacements to $k$ to maintain the subtlety of the perturbations. The optimization process for character replacement also follows a two-step method. Firstly, we calculate the importance score for each character, similar to the approach described in Eq.4. The primary difference is that we consider the removal of individual characters, not words. The top-$k$ characters with the highest importance scores are earmarked for replacement. Secondly, we employ the same greedy search strategy used in synonym replacement to identify the optimal replacements for the selected characters. Note that our character set encompasses uppercase and lowercase letters, digits, punctuation marks, and whitespace, in line with the sets in (Kim et al., 2016; Ebrahimi et al., 2017). The detailed algorithm and its implementation are shown in Algorithm 2 in Appendix  A.1.
Adversarial Suffix. The concept of an adversarial suffix, referred to as adding additional tokens at the end of the original text, has shown considerable effectiveness in misleading LLMs (Zou et al., 2023). Thus, in addition to synonym and character replacement, we also adapt this perturbation to evaluate the token-level vulnerability of the ICL process. To ensure imperceptible to humans, we restrict the number of additional tokens when adapting to our framework. For a given text $x$ that can be tokenized into a sequence of tokens $x=[t_{1},...,t_{n}]$, we define $\delta(x)$ as $[t_{1},...,t_{n},t^{\prime}_{1},...,t^{\prime}_{k}]$ where $t^{\prime}_{1},...,t^{\prime}_{k}$ are adversarial suffices. Our goal is to identify the optimal suffixes that maximize the objective in Eq 3. We also employ a greedy search approach, involving iteratively selecting each suffix token from $t^{\prime}_{1}$ to $t^{\prime}_{k}$ one by one which results in the maximum increase in the loss. The detailed implementation and optimization process for this approach is elaborated in Algorithm 3 in Appendix  A.1.
4 Experiment
We conduct extensive experiments to validate the effectiveness of the proposed framework ICLPoison, particularly with three perturbations introduced in Section 3.3.
4.1 Experiments setting
Datasets. We conduct experiments on different types of classification tasks. Stanford Sentiment Treebank (SST2) dataset from the GLUE benchmark (Wang et al., 2019) is a sentiment analysis dataset consisting of sentences from movie reviews and human annotations of their sentiment in 2 classes (positive/negative); Corpus of Linguistic Acceptability (Cola) dataset from GLUE is a linguistic analysis dataset consisting of English acceptability judgments collected from linguistic books, labeled with “acceptable” or “unacceptable”; Emo dataset (Wang et al., 2023) focuses on emotion classification consisting of Twitter messages labeled in 4 classes; AG’s new (AG) corpus (Zhang et al., 2015) is a topic classification dataset gathered from news sources and labeled in 4 classes (World, Sports, Business, and Science/Technology); and Poem Sentiment (Poem) (Sheng & Uthus, 2020) is a sentiment analysis dataset of poem verses from Project Gutenberg, classified into 3 classes.
Models. In this work, we consider both open-source models including Llama2-7B (Touvron et al., 2023), Pythia (2.8B, 6.9B) (Biderman et al., 2023), Falcon-7B (Almazrouei et al., 2023), GPT-J-6B (Wang & Komatsuzaki, 2021), MPT-7B (Team, 2023), and API-only models GPT-3.5 and GPT-4 (Brown et al., 2020).
Baselines. Since we are the first to study poisoning attacks in ICL, we compare our methods with clean ICL and random label flip (Min et al., 2022). For the baseline random label flip, we replace the true label of the example with a random label uniformly selected from the label space.
Metrics. Our main focus is on the ICL accuracy. For every dataset, we generate perturbations for examples in the training data and randomly select examples from it to conduct ICL prediction for every sample in the test data. We repeat for 5 runs and report the average ICL accuracy. We also include perplexity scores, which are defined as the average negative log-likelihood of each of the tokens appearing, showing whether the perturbations are imperceptible or not.
[…]
Table 1: Main results for attacking LLMs. Average ICL accuracy on clean data and poisoned data (5 independent runs) as well as standard error are reported (in percentage), where lower accuracy represents a stronger poisoning effect. The lowest accuracy for each row is highlighted in blue.
Model Dataset Clean Random label Synonym Character Adv suffix
Cola 55.2$\pm$1.8 49.3$\pm$2.1 10.4$\pm$2.1 17.6$\pm$1.1 13.8$\pm$1.3
SST2 82.8$\pm$1.4 79.4$\pm$1.9 19.4$\pm$1.8 23.8$\pm$1.9 22.7$\pm$1.6
Emo 70.3$\pm$2.3 39.4$\pm$2.2 12.5$\pm$1.1 14.7$\pm$1.4 10.4$\pm$1.5
Poem 56.2$\pm$1.8 43.1$\pm$2.3 12.3$\pm$1.5 17.9$\pm$1.2 13.8$\pm$1.1
Pythia-6.9B AG 66.5$\pm$2.1 47.9$\pm$1.7 13.8$\pm$1.3 17.3$\pm$1.5 12.9$\pm$1.7
Cola 63.8$\pm$1.9 55.5$\pm$2.0 15.3$\pm$1.7 22.7$\pm$2.1 13.6$\pm$1.4
SST2 88.6$\pm$1.5 82.1$\pm$3.2 18.5$\pm$2.0 26.8$\pm$1.7 20.4$\pm$1.7
Emo 73.1$\pm$1.3 43.6$\pm$1.9 11.9$\pm$1.8 17.5$\pm$1.4 12.7$\pm$1.3
Poem 62.9$\pm$1.8 51.4$\pm$2.3 18.1$\pm$1.9 23.3$\pm$1.6 17.2$\pm$1.1
Llama2-7B AG 73.2$\pm$2.0 57$\pm$2.6 13.6$\pm$2.2 19.4$\pm$1.3 11.9$\pm$1.2
Cola 65.2$\pm$1.5 44.8$\pm$1.7 12.7$\pm$1.9 16.5$\pm$1.7 10.8$\pm$1.4
SST2 83.8$\pm$2.5 83.1$\pm$2.5 20.1$\pm$1.6 25.8$\pm$1.3 22.7$\pm$1.7
Emo 61.1$\pm$1.7 52.6$\pm$1.9 10.8$\pm$1.5 14.1$\pm$1.9 9.9$\pm$1.1
Poem 55.2$\pm$1.4 42.9$\pm$1.5 10.5$\pm$1.9 17.3$\pm$1.5 13.6$\pm$1.3
Falcon-7B AG 75.2$\pm$1.8 50.8$\pm$1.3 11.2$\pm$2.3 14.9$\pm$1.7 12.8$\pm$1.2
Cola 57.8$\pm$1.3 49.1$\pm$2.5 13.7$\pm$1.7 17.2$\pm$1.8 11.8$\pm$0.9
SST2 85.4$\pm$1.6 82.8$\pm$2.1 14.8$\pm$2.0 18.9$\pm$1.5 11.4$\pm$1.1
Emo 58.7$\pm$1.1 46.2$\pm$1.7 11.7$\pm$1.8 13.8$\pm$1.3 9.6$\pm$0.7
Poem 57.6$\pm$1.5 46.7$\pm$1.4 12.6$\pm$2.4 14.2$\pm$2.2 10.3$\pm$1.3
GPT-J-6B AG 63.2$\pm$1.7 53.4$\pm$1.9 11.9$\pm$1.5 16.8$\pm$1.8 12.5$\pm$1.1
Cola 75.6$\pm$0.7 76.3$\pm$0.6 58.1$\pm$0.5 62.6$\pm$0.4 59.7$\pm$0.4
SST2 93.8$\pm$0.3 89.7$\pm$0.5 76.8$\pm$0.2 78.3$\pm$0.5 74.2$\pm$0.9
Emo 73.8$\pm$0.5 72.4$\pm$0.8 65.4$\pm$0.4 63.1$\pm$0.7 61.3$\pm$0.5
Poem 51.4$\pm$0.9 53.3$\pm$0.6 39.7$\pm$0.6 45.2$\pm$0.6 43.9$\pm$0.4
GPT-3.5-turbo AG 85.6$\pm$0.3 80.7$\pm$0.4 76.2$\pm$0.5 73.8$\pm$0.2 69.4$\pm$0.7
Cola 85.8$\pm$0.2 82.1$\pm$0.3 73.1$\pm$0.5 75.8$\pm$0.3 69.6$\pm$0.4
SST2 95.1$\pm$0.4 92.5$\pm$0.5 81.5$\pm$0.2 86.1$\pm$0.2 82.3$\pm$0.5
Emo 84.9$\pm$0.1 81.7$\pm$0.2 80.9$\pm$0.6 78.1$\pm$0.5 78.3$\pm$0.4
Poem 72.4$\pm$0.2 63.8$\pm$0.7 56.7$\pm$0.9 60.9$\pm$0.7 57.1$\pm$0.3
GPT-4 AG 90.4$\pm$0.3 87.3$\pm$0.3 83.2$\pm$0.5 83.1$\pm$0.4 84.7$\pm$0.5
4.2 Main results
In this subsection, we apply our framework to various LLMs and datasets to evaluate the vulnerability of ICL.
Attacking open-source models. Our study initially examines open-source models. For each, we craft poisoned samples utilizing the model’s own architecture and subsequently assess the ICL accuracy. Note that we will discuss the transferability i.e. examining the performance of poisoned samples crafted using Llama2-7B when tested on different models in Section 4.3. We begin by altering the entire training set to create these samples. Further studies involving only a subset of poisoned data are presented in Section 4.4. Partial results are shown in Table 1 and full results can be found in Table 5 in Appendix A.2. In the table, a lower accuracy indicates a stronger poisoning effect and the lowest performance is highlighted. It is obvious that ICL performs well on clean data, especially for the SST2 dataset and Llama2-7B model, achieving more than 88% accuracy. While random labels do interrupt the ICL process, causing a decrease in performance, LLMs are still capable of learning effectively from the demonstrations with the dropping of ICL accuracy less than 7%, which is aligned with observations by (Min et al., 2022). In contrast, our ICLPoison framework significantly reduces ICL accuracy, achieving drops to below 10% for some models and datasets such as GPT-J-6B with the Emo dataset. Notably, ICLPoison reveals that ICL is vulnerable to different levels of data poisoning attacks, ranging from subtle character-level changes to word-level manipulations. Among all three variants of ICLPoison, synonym replacement and adversarial suffix perturbations exhibit more severe threats than character replacement in reducing ICL accuracy, likely because they introduce more changes within the same perturbation budget. We present some poisoned examples from three methods in Appendix A.3 for human evaluation. We observe that the poisoning effect differs across models and datasets. For instance, Llama2-7B shows heightened sensitivity to poisoning on the Emo dataset, whereas Falcon-7B is particularly vulnerable to attacks on the Poem Sentiment dataset. These variations may stem from differences in model architectures and the inherent complexity of the datasets.
Figure 2: Experimental results of transferring poisoned data from Llama2-7B to other models. The Y-axis represents the ICL accuracy (a smaller value represents a stronger poisoning effect), while the X-axis denotes different models.
Attacking API-only models. For API-only models like GPT-3.5-turbo and GPT-4, we lack direct access to their internal model representations. Therefore, we employ Llama2-7B as a surrogate to generate poisoned samples and assess the ICL accuracy using the provided APIs. The outcomes, detailed in Table 1, reveal that our approach using Llama2-7B effectively reduces the ICL accuracy of these cutting-edge models by about 10% for both GPT-3.5 and GPT-4. This not only validates the effectiveness of our method but also confirms its utility in real-world applications with advanced LLMs. Additionally, we observe that compromising such models poses greater challenges than open-source models, potentially due to their larger scale and the use of surrogate models (because of the black-box nature). Furthermore, these models display varying degrees of vulnerability to different perturbation intensities. Notably, GPT-4 exhibits particular susceptibility to character replacement, suggesting a heightened sensitivity to minor textual variations.
4.3 Transferbility
In this subsection, we consider a practical scenario when the attacker has no access to the victim LLM, and leverages a surrogate model to generate poisoned samples. Specifically, we adopt the Llama2-7B model in this study. This setup allows us to evaluate the transferability of our poisoning approach across different models, including black-box models such as GPT-3.5 and GPT-4. The initial results for the transferability study are presented in Figure 2, in which we focus on the SST2 dataset. For a more comprehensive analysis covering all five datasets used in our study, we direct readers to Appendix A.2.
From Figure 2, we observe a notable trend: while the efficacy of the poisoning attack diminishes when moving from the surrogate (Llama2-7B) to the victim models, the impact remains significant. The poisoned examples generated by our ICLPoison framework – in all three of its variants – substantially jeopardize the performance of ICL, leading to over a 30% decrease in accuracy for open-source models. Furthermore, our analysis sheds light on the different transferability of various perturbations. Synonym replacement and adversarial suffixes, in particular, demonstrate a stronger poisoning effect compared to character replacement. This disparity could be attributed to the more pronounced influence these methods exert on the surrogate model. We also note the influence of model size on susceptibility to poisoning. Models smaller in size than the surrogate, such as Pythia-2.8B and GPT-J-6B, appear to be more vulnerable to these poisoning examples, while larger models exhibit a degree of resistance. This pattern suggests that the effectiveness of our approach may be modulated by the size and complexity of the target model, offering valuable insights for future research in trustworthy language models.
4.4 Partial poisoning
To consider practical scenarios where attackers can only alter a part of the data, our experiments involve perturbing a random subset of the training dataset. We allocate varying proportions (10%, 20%, 50%, and 100%) of the training data as accessible for manipulation in the Llama2-7B model and GLUE-SST2 dataset. The results, shown in Figure 3, reveal that a lower mixing rate reduces the overall poisoning effect. Nonetheless, even at a 10% rate, we observe a significant decrease in ICL performance by over 7%, and a 15% decrease at 20%, underscoring the efficacy of our ICLPoison framework.
Figure 3: Results for partially poisoned ICL. The X-axis represents the ratio of poisoned data while Y-axis represents the ICL accuracy.
These findings highlight the vulnerability of ICL to subtle data poisoning attacks, where even a limited number of malicious inputs can significantly disrupt the ICL process. Additionally, the results show that perturbation strategies like synonym replacement and adversarial suffix, which introduce more pronounced textual changes within a fixed poisoning budget, more severely affect ICL performance compared to character replacement. This indicates that broader, coarse-grained perturbations are generally more disruptive than finer, more subtle ones.
Table 2: Average perplexity scores for clean and poisoned data with standard error reported. Focus on model Llama2-7B.
Dataset Clean Synonym Character Adv suffix
Cola 4.66$\pm$1.06 5.22$\pm$1.00 7.37$\pm$0.89 9.58$\pm$1.08
SST2 5.48$\pm$1.16 6.66$\pm$0.73 7.45$\pm$0.70 7.75$\pm$1.13
Emo 5.02$\pm$1.02 5.79$\pm$0.67 7.27$\pm$0.79 7.31$\pm$0.75
Poem 5.39$\pm$0.85 6.32$\pm$0.62 8.64$\pm$0.49 9.27$\pm$1.09
AG 2.37$\pm$0.41 3.12$\pm$0.29 3.84$\pm$0.50 3.68$\pm$0.35
4.5 Potential defenses
To evaluate the robustness of our framework, we applied two representative defenses from (Jain et al., 2023): detection-based defense perplexity filter and preprocessing defense paraphrasing.
Perplexity filter. The perplexity score of a text is referred to as the average negative log-likelihood of each of the tokens appearing. The perturbation added to the original text may cause grammar mistakes, logic problems, and a reduction in fluency. This will increase the perplexity, and thus the corresponding text could be more detectable (Jain et al., 2023). We report the perplexity scores of generated poisoned data, and a higher value indicates a higher probability of being detected and lower robustness. Table 2 shows the perplexity scores for poisoned data across various models and datasets, calculated as described in Section 4.1 of (Jain et al., 2023). Among the perturbations, synonym replacement yields relatively lower perplexity most close to the clean perplexity, whereas adversarial suffixes generally result in the highest scores. This suggests that synonym replacement is less detectable than the adversarial suffix.
Paraphrasing. Paraphrasing is a typical preprocessing method that utilizes a language model to rewrite the input text. This process aims to maintain original meanings while removing adversarial perturbations, making it an effective defensive strategy. In this part, we leverage the cutting-edge GPT-4 model to paraphrase the poisoned data and evaluate ICL performance on paraphrased inputs. As shown in Table 3, paraphrasing successfully neutralizes adversarial suffixes while largely preserving the impact of synonym replacements. This outcome is anticipated since adversarial suffixes typically introduce irrelevant content, whereas synonym replacements maintain the text’s semantic integrity.
Table 3: The poisoned data is paraphrased by GPT-4, and ICL accuracy is reported. Original results are included in brackets.
Datasets Clean Random label Synonym Character Adv suffix
Cola 65.2(63.8) 53.9(55.5) 36.5(15.3) 50.6(22.7) 58.5(13.6)
SST2 83.1(88.6) 85.4(82.1) 52.1(18.5) 60.2(26.8) 80.2(20.4)
Emo 75.5(73.1) 48.2(43.6) 40.7(14.9) 48.3(17.5) 66.8(12.7)
Poem 63.7(62.9) 52.1(51.4) 34.3(18.1) 43.7(23.3) 55.2(17.2)
AG 70.6(73.2) 55.7(57) 38.2(13.6) 47.2(19.4) 64.3(11.9)
The above findings make it evident that the implemented defenses can enhance the robustness of ICL against data poisoning to some extent. However, the degree of improvement is contingent upon the type of perturbation. For instance, token-level perturbations, such as adversarial suffixes, are more readily addressed by these defenses, whereas word-level perturbations like synonym replacements pose a greater threat due to their subtlety and ability to alter semantic meaning. Character-level perturbations fall between these two in terms of detectability and potential impact, with their severity likely to increase with a higher attack budget. Our results highlight the necessity to develop a robust ICL.
4.6 Ablation studies
Under the assumption that attackers lack knowledge of the ICL process details, such as the templates and the number of examples used, we perform ablation studies to understand the impact of these settings. Specifically, we evaluate three different templates: F1, which is used in generating poisoned data; F2, formatted as ’input-output’; and F3, structured as ’Q:input, A:output’. Additionally, we explore the effects of using 3, 5, and 7 examples in ICL predictions. This ablation study concentrates on the Llama2-7B model and the GLUE-SST2 dataset, with comprehensive results provided in Appendix A.2. The results for different templates and varying numbers of examples are detailed in Table 4 and Figure 4, respectively. The results demonstrate that our framework is effectively adaptable to various templates and different numbers of demonstrations, highlighting its potentials in practice.
Table 4: Experiments with different templates: F1, F2, F3 on model Llama2-7B and dataset GLUE-SST2.
F1 F2 F3
Clean 88.6 92.5 90.3
Random label 82.1 84.7 79.2
Synonym 18.5 17.9 18.2
Character 26.8 30.6 28.5
Adv suffix 20.4 21.7 19.3
Figure 4: Experiments with different numbers of examples.
5 Related Works
5.1 In-context Learning
In-context learning (ICL), since its introduction by (Brown et al., 2020), has been notable for efficiently tackling tasks with just a few examples, bypassing the need to adjust the model’s parameters (Dong et al., 2022; Liu et al., 2021; Sun et al., 2022). Generally speaking, ICL allows LLMs to make predictions based on a few examples formatted in a template rather than changing model parameters. Existing works have shown that the performance of ICL relies on the quality of demonstrations, including the selection of examples (Liu et al., 2021; Sorensen et al., 2022; Wang et al., 2023), the order of examples (Lu et al., 2021; Min et al., 2022), the template used to format demonstrations (Liu et al., 2023a; Wei et al., 2022). The mechanisms behind ICL’s success have been the subject of much interest. One line of work reveals that ICL learns latent concepts and conducts implicit Bayesian inference (Xie et al., 2021). Follow-up works (Hendel et al., 2023; Liu et al., 2023b) take a further step and point out that hidden states in LLMs encode these concepts. Another line of work focuses on the implicit learning mechanism of ICL. Garg et al. (2022) showed that Transformers (Vaswani et al., 2017) can encode effective learning algorithms to learn unseen linear functions according to demonstration samples. Von Oswald et al. (2023); Dai et al. (2022); Akyürek et al. (2022) connect Transformers with gradient descent and claim that LLMs can implicitly perform gradient descent on examples. Bai et al. (2023) further extend the ability of implicit gradient descent to various statistical abilities including pre-ICL testing and post-ICL validation. Our work builds on the foundation of the first line, and leverages the understanding of latent concepts and hidden states to conduct effective data poisoning attacks.
5.2 Data Poisoning
Data poisoning attacks (Biggio et al., 2012; Steinhardt et al., 2017) traditionally occur during the data collection phase of machine learning model training, where the training data is tampered with to induce malicious behaviors in the resulting models. These behaviors can range from degraded performance on testing data (Steinhardt et al., 2017; Huang et al., 2021) to the misclassification of specific test samples (Shafahi et al., 2018; Zhu et al., 2019) and the implantation of backdoors. In conventional models, such attacks typically exploit the training objectives (Steinhardt et al., 2017; He et al., 2023). However, applying data poisoning to in-context learning (ICL) aims to undermine the ICL’s overall effectiveness, presenting unique challenges due to its implicit learning mechanism.
6 Conclusion
In this study, we introduce ICLPoison, a novel framework devised to assess the vulnerability of in-context learning (ICL) in the face of data poisoning attacks. We use the dynamics of hidden states in Large Language Models (LLMs) to craft our attack objectives. Furthermore, we implement our framework through three distinct and practical algorithms, each employing a different method of discrete perturbation. Our research exposes previously unidentified susceptibilities of the ICL process to data poisoning. This discovery emphasizes the urgent need for enhancing the robustness of ICL implementations.
Broader Impact
This paper introduces a new framework to assess how in-context learning in large language models (LLMs) is susceptible to data poisoning attacks. It highlights the vulnerabilities in in-context learning, urging the development of more robust implementations. Additionally, the framework could aid in creating stronger attacks. Our findings underscore the pressing need to enhance LLMs’ security to protect against such vulnerabilities.
Appendix A Appendix
A.1 Details of Algorithms
In this section, we present detailed algorithms, including synonym replacement in Algorithm 1, character replacement in Algorithm 2 and adversarial suffix in Algorithm 3.
Algorithm 1. Algorithm 1 describes the whole process of conducting ICLPoison with synonym replacement. For each example $x^{p}_{i,t}$ in the accessible prompting set $D^{p}_{t}$, it first selects words to be replaced based on an importance score (step (1)-(3)): the importance score for every word in $x^{p}_{i,t}$ is computed via Eq.4 which is the distortion between the text before and after removing the word (step (2)); then the score for every word is sorted in descending order and $k$ words with largest scores are chosen (step (3)). Secondly (step (4)-(8)), we greedily search for the optimal replacement for selected words within their synonyms: first extract $m$ synonyms based on cosine similarity of GloVe embeddings (step (4)); then each selected word is replaced with its synonyms and evaluates the distortion with the original text via Eq. 2 (step (5)-(6)); the synonym causing the largest distortion is chosen as the final replacement.
Algorithm 2. Algorithm 2 describes the whole process of conducting ICLPoison with character replacement. For each example $x^{p}_{i,t}$ in the accessible prompting set $D^{p}_{t}$, it first selects characters to be replaced based on an importance score (step (1)-(3)): the importance score for every character in $x^{p}_{i,t}$ is computed via Eq.4 which is the distortion between the text before and after removing the character (step (2)); then the score for every character is sorted in descending order and $k$ words with largest scores are chosen (step (3)). Secondly (step (4)-(7)), we greedily search for the optimal replacement for selected characters within the character set $C$: each selected word is replaced with characters inside $C$ and evaluates the distortion with the original text via Eq. 2 (step (4)-(5)); the character causing the largest distortion is chosen as the final replacement.
Algorithm 3. Algorithm 3 describes the whole process of conducting ICLPoison with adversarial suffix. For each example $x^{p}_{i,t}$ in the accessible prompting set $D^{p}_{t}$, it first randomly initializes the $k$ suffices (step (2)). We greedily search for the optimal token for each suffix within the vocabulary set $V$: each suffix is replaced with tokens inside $V$ and evaluates the distortion with the original text via Eq. 2 (step (3)-(4)); the character causing the largest distortion is chosen as the final replacement.
Algorithm 1 ICLPoison + Synonym replacement
Input Clean prompting set $D_{t}^{p}=\{(x^{p}_{i,t},y^{p}_{i,t})\}_{i=1}^{N}$, surrogate model $f$ consisting of $L$ layers, attacking budget $k$, number of synonyms $m$.
Output Poisoned prompting set $\{(\delta_{i}(x^{p}_{i,t}),y^{p}_{i,t})\}_{i=1}^{N}$
  for $i=1,...,N$ do
     Step 1: Select words to replace with importance scores
     (1) Decompose input text $x^{p}_{i,t}$ into a sequence of words $[w_{1},...,w_{n}]$
     (2) Compute importance score $I_{w_{j}}$ for each word $w_{j}$ with Eq. 4
     (3) Sort scores in descending order $I_{w_{(1)}}\geq I_{w_{(2)}}\geq\cdot\geq I_{w_{(n)}}$ and select top-$k$ words: $w_{(1)},...,w_{(k)}$
     Step 2: Select optimal synonyms for each selected word
     for $w\in[w_{(1)},...,w_{(k)}]$ do
        (4) Obtain top-$m$ synonyms $[s_{(1)},...,s_{(m)}]$ with highest cosine similarity with $w$ based on GloVe word embeddins.
        for $s\in[s_{(1)},...,s_{(m)}]$ do
           (5) Replace $w$ with $s$ obtaining $x^{\prime}_{w}=[w_{1},...,s,...,w_{n}]$
           (6) Evaluate the distortion of hidden states after replacement with Eq.2: $\mathcal{L}_{d}(H(x^{p}_{i,t},f),H(x^{\prime}_{w},f))$
        end for
        (7) Select the synonym causing the largest distortion to replace $w$.
     end for
     (8) Obtain perturbed input $\delta_{i}(x^{p}_{i,t})$
  end for
  Return poisoned prompting set $\{(\delta_{i}(x^{p}_{i,t}),y^{p}_{i,t})\}_{i=1}^{N}$
Algorithm 2 ICLPoison + Character replacement
Input Clean prompting set $D_{t}^{p}=\{(x^{p}_{i,t},y^{p}_{i,t})\}_{i=1}^{N}$, surrogate model $f$ consisting of $L$ layers, attacking budget $k$, character set $C$.
Output Poisoned prompting set $\{(\delta_{i}(x^{p}_{i,t}),y^{p}_{i,t})\}_{i=1}^{N}$
  for $i=1,...,N$ do
     Step 1: Select characters to replace with importance scores
     (1) Decompose input text $x^{p}_{i,t}$ into a sequence of characters $[c_{1},...,c_{n}]$
     (2) Compute importance score $I_{c_{j}}$ for each word $c_{j}$ with Eq. 4
     (3) Sort scores in descending order $I_{c_{(1)}}\geq I_{c_{(2)}}\geq\cdot\geq I_{w_{(n)}}$ and select top-$k$ words: $c_{(1)},...,c_{(k)}$
     Step 2: Select optimal character for each selected character from the whole character set.
     for $c\in[c_{(1)},...,c_{(k)}]$ do
        for $c^{\prime}\in C$ do
           (4) Replace $c$ with $c^{\prime}$ obtaining $x^{\prime}_{c}=[c_{1},...,c^{\prime},...,c_{n}]$
           (5) Evaluate the distortion of hidden states after replacement with Eq.2: $\mathcal{L}_{d}(H(x^{p}_{i,t},f),H(x^{\prime}_{c},f))$
        end for
        (6) Select the character causing the largest distortion to replace $c$.
     end for
     (7) Obtain perturbed input $\delta_{i}(x^{p}_{i,t})$
  end for
  Return poisoned prompting set $\{(\delta_{i}(x^{p}_{i,t}),y^{p}_{i,t})\}_{i=1}^{N}$
Algorithm 3 ICLPoison + Adversarial suffix
Input Clean prompting set $D_{t}^{p}=\{(x^{p}_{i,t},y^{p}_{i,t})\}_{i=1}^{N}$, surrogate model $f$ consisting of $L$ layers, attacking budget $k$, token vocabulary $V$.
Output Poisoned prompting set $\{(\delta_{i}(x^{p}_{i,t}),y^{p}_{i,t})\}_{i=1}^{N}$
  for $i=1,...,N$ do
     (1) Tokenize text $x^{p}_{i,t}$ into sequence of tokens $[t_{1},...,t_{n}]$
     (2) Random initialize the adversarial suffix and concatenate with the original text: $\delta(x^{p}_{i,t})=[t_{1},...,t_{n},t_{1}^{\prime},...,t_{k}^{\prime}]$
     for $j\in[k]$ do
        for $v\in V$ do
           (3) Replace $t_{j}^{\prime}$ with $v$ obtaining $x^{\prime}_{t}=[t_{1},...,t_{n},t^{\prime}_{1},...,v,...,t^{\prime}_{k}]$
           (4) Evaluate the distortion of hidden states after replacement with Eq.2: $\mathcal{L}_{d}(H(x^{p}_{i,t},f),H(x^{\prime}_{t},f))$
        end for
        (5) Select the token causing the largest distortion to replace $t_{j}^{\prime}$.
     end for
     (6) Obtain perturbed input $\delta_{i}(x^{p}_{i,t})$
  end for
  Return poisoned prompting set $\{(\delta_{i}(x^{p}_{i,t}),y^{p}_{i,t})\}_{i=1}^{N}$
A.2 Additional Experiments
In this section, we present additional experimental results, including full results on attacking open-source models in Table 5, full results of transferability in Table 7, full results of perplexity scores in Table 8, full results on various templates in Table 6 and the number of examples in Table 9.
Attack open-source models. In Table 5, we include more results on additional models such as Pythia-2.8B and MPT-7B. Our observation is consistent with the analysis in Section 4.2.
Transferbility. In Table 7, results on all 5 datasets are presented, and we notice that the transferability of three perturbations varies. This may be because of the capacity of models and the complexity of datasets. A detailed investigation can be an interesting future direction.
Perplexity scores. Table 8 covers perplexity scores on various datasets and models. It is obvious that synonym replacement is more stealthy than the other 2 methods.
Impact of templates. We test 3 different templates on various datasets and models. Our results in Table 6 reveal that our poisoned examples remain effective across templates.
Impact of the number of examples. Our results about different numbers of examples in Table 9 show that more examples can improve ICL performance, while also leading to easier manipulation and stronger poisoning effect.
Table 5: Full results on attacking open-source models
Model Dataset Clean Random label Synonym Character Adv suffix
Cola 64.1$\pm$1.6 59.4$\pm$1.6 12.3$\pm$1.2 17.6$\pm$1.4 14.2$\pm$1.3
SST2 76.5$\pm$1.5 70.2$\pm$1.7 17.5$\pm$1.1 24.3$\pm$1.2 18.1$\pm$2.1
Emo 67.2$\pm$1.3 48.1$\pm$2.8 10.9$\pm$1.8 15.7$\pm$1.7 12.3$\pm$1.7
Poem 57.1$\pm$1.7 31.8$\pm$1.9 10.5$\pm$1.6 16.4$\pm$1.6 9.7$\pm$1.2
Pythia-2.8B AG 59.4$\pm$1.1 46.5$\pm$1.7 15.7$\pm$1.0 20.3$\pm$1.3 14.6$\pm$1.4
Cola 55.2$\pm$1.8 49.3$\pm$2.1 10.4$\pm$2.1 17.6$\pm$1.1 13.8$\pm$1.3
SST2 82.8$\pm$1.4 79.4$\pm$1.9 19.4$\pm$1.8 23.8$\pm$1.9 22.7$\pm$1.6
Emo 70.3$\pm$2.3 39.4$\pm$2.2 12.5$\pm$1.1 14.7$\pm$1.4 10.4$\pm$1.5
Poem 56.2$\pm$1.8 43.1$\pm$2.3 12.3$\pm$1.5 17.9$\pm$1.2 13.8$\pm$1.1
Pythia-6.9B AG 66.5$\pm$2.1 47.9$\pm$1.7 13.8$\pm$1.3 17.3$\pm$1.5 12.9$\pm$1.7
Cola 63.8$\pm$1.9 55.5$\pm$2.0 15.3$\pm$1.7 22.7$\pm$2.1 13.6$\pm$1.4
SST2 88.6$\pm$1.5 82.1$\pm$3.2 18.5$\pm$2.0 26.8$\pm$1.7 20.4$\pm$1.7
Emo 73.1$\pm$1.3 43.6$\pm$1.9 11.9$\pm$1.8 17.5$\pm$1.4 12.7$\pm$1.3
Poem 62.9$\pm$1.8 51.4$\pm$2.3 18.1$\pm$1.9 23.3$\pm$1.6 17.2$\pm$1.1
Llama2-7B AG 73.2$\pm$2.0 57$\pm$2.6 13.6$\pm$2.2 19.4$\pm$1.3 11.9$\pm$1.2
Cola 65.2$\pm$1.5 44.8$\pm$1.7 12.7$\pm$1.9 16.5$\pm$1.7 10.8$\pm$1.4
SST2 83.8$\pm$2.5 83.1$\pm$2.5 20.1$\pm$1.6 25.8$\pm$1.3 22.7$\pm$1.7
Emo 61.1$\pm$1.7 52.6$\pm$1.9 10.8$\pm$1.5 14.1$\pm$1.9 9.9$\pm$1.1
Poem 55.2$\pm$1.4 42.9$\pm$1.5 10.5$\pm$1.9 17.3$\pm$1.5 13.6$\pm$1.3
Falcon-7B AG 75.2$\pm$1.8 50.8$\pm$1.3 11.2$\pm$2.3 14.9$\pm$1.7 12.8$\pm$1.2
Cola 57.8$\pm$1.3 49.1$\pm$2.5 13.7$\pm$1.7 17.2$\pm$1.8 11.8$\pm$0.9
SST2 85.4$\pm$1.6 82.8$\pm$2.1 14.8$\pm$2.0 18.9$\pm$1.5 11.4$\pm$1.1
Emo 58.7$\pm$1.1 46.2$\pm$1.7 11.7$\pm$1.8 13.8$\pm$1.3 9.6$\pm$0.7
Poem 57.6$\pm$1.5 46.7$\pm$1.4 12.6$\pm$2.4 14.2$\pm$2.2 10.3$\pm$1.3
GPT-J-6B AG 63.2$\pm$1.7 53.4$\pm$1.9 11.9$\pm$1.5 16.8$\pm$1.8 12.5$\pm$1.1
Cola 53.4$\pm$1.2 45.3$\pm$1.2 15.6$\pm$1.6 17.4$\pm$1.9 14.1$\pm$1.4
SST2 89$\pm$1.5 82.9$\pm$2.3 20.4$\pm$1.9 25.6$\pm$2.5 19.8$\pm$1.3
Emo 59.7$\pm$1.3 41.8$\pm$1.7 9.6$\pm$1.5 11.5$\pm$1.6 10.4$\pm$0.8
Poem 69$\pm$1.8 56.2$\pm$2.5 14.9$\pm$1.4 16.3$\pm$1.5 12.7$\pm$1.2
MPT-7B AG 70.6$\pm$1.6 55.3$\pm$1.9 13.9$\pm$1.7 17.1$\pm$1.9 15.2$\pm$1.6
Table 6: Evaluating data poisoning attacks on different ICL templates. F1, F2, F3 denote 3 different templates, and ICL accuracy on various dataset is reported.
Dataset ICL format Clean Random label Synonym Character Adv suffix
Cola F1 63.8 55.5 15.3 22.7 13.6
F2 55.1 47.5 14.6 20.9 13.9
F3 59.5 53.3 13.8 19.4 12.5
SST2 F1 88.6 82.1 18.5 26.8 20.4
F2 92.5 84.7 17.9 30.6 21.7
F3 90.3 79.2 18.2 28.5 19.3
Emo F1 73.1 43.6 14.9 17.5 12.7
F2 66.7 37.6 11.6 12.9 9.2
F3 69.1 39.8 12.8 15.6 11.3
Poem F1 62.9 51.4 18.1 23.3 17.2
F2 57.1 45.2 13.7 20.1 12.8
F3 61.9 49.5 16.3 21.5 13.7
AG F1 73.2 57.6 13.6 19.4 11.9
F2 68.6 54.6 11.7 18.2 10.3
F3 75.8 67.2 14.1 19.3 12.6
Table 7: Full results for testing poisoned examples generated by Llama2-7B on other models.
Dataset Clean Random label Synonym Character Adv suffix
Cola 64.1 59.4 31.9 34.0 34.8
SST2 76.5 70.2 38.5 33.9 36.6
Emo 67.2 48.1 26.3 32.8 30.3
Poem 57.1 31.8 33.3 27.1 24.3
Pythia-2.8B AG 59.4 46.5 31.7 31.6 35.6
Cola 55.2 49.3 26.7 34.7 34.2
SST2 82.8 79.4 39.5 41.0 41.3
Emo 70.3 39.4 18.9 24.1 20.9
Poem 56.2 43.1 29.0 28.1 26.2
Pythia-6.9B AG 66.5 47.9 32.1 33.9 27.2
Cola 65.2 44.8 23.4 24.8 25.0
SST2 83.8 83.1 37.7 35.8 34.7
Emo 61.1 52.6 28.6 30.5 27.1
Poem 55.2 42.9 25.1 27.3 25.6
Falcon-7B AG 75.2 50.8 24.7 24.8 24.2
Cola 57.8 49.1 29.7 28.5 29.1
SST2 85.4 82.8 31.7 31.8 30.2
Emo 58.7 46.2 22.3 24.1 19.3
Poem 57.6 46.7 26.7 28.6 23.3
GPT-J-6B AG 63.2 53.4 29.4 29.5 29.8
Cola 53.4 45.3 22.6 23.2 19.3
SST2 89 82.9 29.8 29.6 28.8
Emo 59.7 41.8 21.1 25.4 25.1
Poem 69 56.2 26.2 24.5 23.5
MPT-7B AG 70.6 55.3 31.0 34.2 27.8
Cola 75.6 76.3 58.1 62.6 59.7
SST2 93.8 89.7 76.8 78.3 74.2
Emo 73.8 72.4 65.4 63.1 61.3
Poem 51.4 53.3 39.7 45.2 43.9
GPT-3.5-turbo AG 85.6 80.7 76.2 73.8 69.4
Cola 85.8 82.1 73.1 75.8 69.6
SST2 95.1 92.5 81.5 86.1 82.3
Emo 84.9 81.7 80.9 78.1 78.3
Poem 72.4 63.8 56.7 60.9 57.1
GPT-4 AG 90.4 87.3 83.2 83.1 84.7
Table 8: Perplexity scores for poisoned texts across different models and datasets. A lower value means more logical and fluent expression, and fewer grammar mistakes, thus is more imperceptible to humans.
Dataset Clean Synonym Character Adv suffix
Pythia-2.8B Cola 4.87 5.15 7.38 8.35
SST2 5.54 6.37 7.45 8.80
Emo 5.46 6.07 7.27 6.81
Poem 5.50 6.86 7.65 8.09
AG 3.26 4.01 4.84 5.98
Pythia-6.9B Cola 4.84 5.43 7.78 7.15
SST2 5.56 5.58 7.93 7.06
Emo 5.41 5.92 7.49 6.45
Poem 5.50 5.71 7.94 7.86
AG 3.14 4.10 5.08 3.91
Llama2-7B Cola 4.66 5.22 7.37 9.58
SST2 5.48 6.66 7.45 7.75
Emo 5.02 5.79 7.27 7.31
Poem 5.39 6.32 8.64 9.27
AG 2.37 3.12 3.84 3.68
Falcon-7B Cola 5.02 5.43 7.57 7.26
SST2 4.76 5.33 6.84 7.12
Emo 5.03 5.40 7.26 6.30
Poem 5.53 6.02 7.80 8.07
AG 2.67 3.49 4.50 3.64
MPT-7B Cola 4.91 5.80 7.73 7.18
SST2 5.29 5.45 6.62 6.88
Emo 5.12 5.47 6.40 6.22
Poem 5.43 5.91 7.94 7.72
AG 2.75 3.53 4.50 3.68
GPT-J-6B Cola 5.01 5.37 7.41 7.7
SST2 5.06 5.35 6.92 7.81
Emo 5.37 5.49 6.40 7.33
Poem 5.43 5.72 8.32 8.02
AG 3.12 4.08 4.06 5.04
Table 9: Full results for different number of examples in the demonstration. Focus on model Llama2-7B.
Dataset num_examples Clean Random label Synonym Character Adv suffix
Cola 3 63.2 59.2 16.5 21.8 12.8
5 63.8 55.5 15.3 22.7 13.6
7 63.1 54.6 14.8 22.5 13.2
SST2 3 84.5 80.2 14.3 29.4 19.6
5 88.6 82.1 18.5 26.8 20.4
7 92.7 86.1 19.3 32.1 21.7
Emo 3 58 35.2 11.5 14.6 7.9
5 73.1 43.6 14.9 17.5 12.7
7 79.2 47.4 11.7 15.8 11.3
Poem 3 61 48.6 16.9 20.4 15.6
5 62.9 51.4 18.1 23.3 17.2
7 66.7 55.2 19.5 22.6 18.3
AG 3 66.9 52.8 12.9 17.5 9.5
5 73.2 57 13.6 19.4 11.9
7 78 60.1 13.2 18.9 11.6
A.3 Poisoned Text Examples
In this section, we provide some poisoned examples in Table 10 for human evaluation. The additional tokens (for adversarial suffix) and substitutions (for synonym and character replacement) are highlighted in red. It is obvious that adversarial suffixes can introduce irrelevant or non-sense content to the original text, thus can be easily detected. On the contrast, synonym and character replacements introduce more subtle changes to the text.
Table 10: Examples of poisoned examples for three type of perturbations.
Adv suffix Singh Hangs Close Despite missing half his fairways and a half-dozen putts inside 12 feet, Vijay Singh grinds out a 1-under 69 to stay within three shots of the lead at the Tour Championship. owing payrollapplyConvert
ETA Bombs Put Radical Basque Party in Tight Spot The armed Basque separatist group ETA showed it was not a spent force with Friday #39;s coordinated attacks at five Madrid petrol stations, but it also further isolated the guerrillas #39; closest political allies, Batasuna. Gould}}\$,attempt
far from the woods where, when the sun has set,Sup clustering#absolute
Few writers and any playwrights meet in Vienna.ausemployedBlamkAut
The tree gradually dropped its fruit to the ground.Anti Africa OH Min==
Synonym Temperatures Dip in Europe, sent(Sending) Price of Oil Up 4.8 The price of coarse(crude) oil surging(surged) 4.8 percent in New York yesterday, the biggest rise in more than five months, on stress(concern) that cold weather in Europe will draw heating oil from the West(United) States.
should(would) i queried(ask) give two qursion(questions) yeah(yes) are you marry(married)
Additional(More) bodies found in Nablus US forces uncovered(discovered) more bodies under the northern city of Mosul on Friday, evident(apparent) victims of an bullying(intimidation) campaign by insurgents against Iraq;(’)s fledgling security forces.
Vast(Huge) Black Holes emerged(Formed) Quickly After Big Bang redOrbit- Incredibly substantial(massive) black holes had fully developed(matured) just a billion years after the birth of the cosmos(universe), according to two separate studies.
can yea(you) help think search(finding) girls which(what) are you state(talking) about i need rather(first) girlfriend
Character kidney implant hope The first human trial of an artificial ’bio’j(-)kidney offers c(a) hopq(e)Y( )of a working implant for patients, say exQ()perts.
oo you are veryF( )lucky haha that s(c)an be said pluL(s) i know to maneuver around the roaX(d)s ’ okbuthl(e)n
TennisG(:) Federer warns rivals Roger Federg(e)r believD(e)s he is now cf(a)pable of winning any tournament in tne S(w)orld.
Zimbabwe curbs rights groups j(Z)imbabwe’s parliami(e)nt passes a controversial bill banning international rights grU(o)ups from i(w)orq(k)ing in the country.
sentence: t)(h)eC( )worthy sucS(c)essor to a bettn(e)r toG(m)orrow

### v3

@@
 Abstract
-In the domain of large language models (LLMs), in-context learning (ICL) has been recognized for its innovative ability to adapt to new tasks, relying on examples rather than retraining or fine-tuning.
-This paper delves into the critical issue of ICL’s susceptibility to data poisoning attacks, an area not yet fully explored.
-We wonder whether ICL is vulnerable, with adversaries capable of manipulating example data to degrade model performance.
-To address this, we introduce ICLPoison, a specialized attacking framework conceived to exploit the learning mechanisms of ICL.
-Our approach uniquely employs discrete text perturbations to strategically influence the hidden states of LLMs during the ICL process.
-We outline three representative strategies to implement attacks under our framework, each rigorously evaluated across a variety of models and tasks.
-Our comprehensive tests, including trials on the sophisticated GPT-4 model, demonstrate that ICL’s performance is significantly compromised under our framework.
-These revelations indicate an urgent need for enhanced defense mechanisms to safeguard the integrity and reliability of LLMs in applications relying on in-context learning.
-Machine Learning, ICML
+In-context learning (ICL) has emerged as a capability of large language models (LLMs), enabling them to adapt to new tasks using provided examples.
+While ICL has demonstrated its strong effectiveness, there is limited understanding of its vulnerability against potential threats.
+This paper examines ICL’s vulnerability to data poisoning attacks.
+We introduce ICLPoison, an attacking method specially designed to exploit ICL’s unique learning mechanisms by identifying discrete text perturbations that influence LLM hidden states.
+We propose three representative attack strategies, evaluated across various models and tasks.
+Our experiments, including those on GPT-4, show that ICL performance can be significantly compromised by these attacks, highlighting the urgent need for improved defense mechanisms to protect LLMs’ integrity and reliability.
+Data Poisoning for In-context Learning
+Pengfei He1thanks: Corresponding to hepengf1@msu.edu, Han Xu2, Yue Xing1, Hui Liu1, Makoto Yamada3, Jiliang Tang1 1Michigan State University  2University of Arizona 3 Okinawa Institute of Science and Technology
 1 Introduction
-In recent years, In-Context Learning (ICL) (Brown et al., 2020; Min et al., 2022) has emerged as an important component of large language models (LLMs).
-Unlike traditional machine learning algorithms that require extensive retraining or fine-tuning to adapt new data (Hoi et al., 2021; Zhang & Yang, 2021; Zhuang et al., 2020), ICL enables LLMs to make predictions based on extra information in the prompt (e.g.
-a few examples related to a specific task), without changing the model parameters.
+In-context learning (ICL) (Brown et al., 2020; Min et al., 2022) has emerged as an important capability of large language models (LLMs).
+Unlike traditional machine learning algorithms that require extensive retraining or fine-tuning to adapt to new tasks (Hoi et al., 2021; Zhang and Yang, 2021; Zhuang et al., 2020), ICL enables LLMs to make predictions based on a few examples related to a specific task, without changing the model parameters.
 For example, consider the task of predicting a person’s nationality and the prompt consists of examples, e.g.
 “Albert Einstein was German; Isaac Newton was English;”, followed by the query “Thomas Edison was”, an LLM such as GPT-4 will predict “American” accurately.
-The efficiency and flexibility of ICL have gained significant attention and revolutionized various real-world applications, ranging from reasoning (e.g.
-chain-of-thought (Wei et al., 2022; Wang et al., 2022)) to retrieval-augmented-generation (RAG) (Lewis et al., 2020).
-The success of ICL depends critically on the examples it utilizes.
-Studies have shown that the ICL performance is sensitive to certain characteristics of demonstrations, e.g., the selection of examples (Wang et al., 2023) and the order of examples in demonstration (Min et al., 2022).
+The efficiency and flexibility of ICL have drawn significant attention and revolutionized various real-world applications, ranging from LLM-based few-shot healthcare agents (Shi et al., 2024) to knowledge tagging in the education domain (Li et al., 2024), where demonstrations are utilized to handle specific tasks.
+Despite the success of ICL, studies have shown that the ICL performance is sensitive to certain characteristics of demonstrations, such as the example selection strategy and the quality of examples (Wang et al., 2023; Min et al., 2022).
 Therefore, it naturally raises a question: Is ICL vulnerable to potential data poisoning attacks?
-In reality, malicious actors may manipulate demonstration data to degrade the model performance.
-For example, adversaries can execute data poisoning attacks by strategically altering the examples in demonstrations used in ICL.
-Such attacks aim to disrupt the learning process, leading to inaccurate or biased predictions.
-In this paper, we aim to answer the above question by delving into data poisoning attacks in ICL and uncovering the vulnerability of ICL against these attacks.
-We consider the standard pipeline of ICL, where examples are randomly selected from a certain data set to tailor ICL for a downstream task.
-In terms of the attack, we assume that the adversary deliberately alters some examples in the data, and his goal is to ensure that the learning process is adversely affected and the overall performance decreases.
-This scenario is both significant and practical.
-For instance, in a RAG system where demonstrations are retrieved from a dataset, the attacker manipulates some data from a specific domain or a task (e.g.
-reviews about a specific brand) to downgrade the quality of the generated content of LLM related to that domain/task.
-To the best of our knowledge, we are the first to investigate the vulnerability of ICL against data poisoning attacks.
-Data poisoning in ICL faces both unique challenges specific to ICL and common obstacles in traditional data poisoning.
-First, in contrast to traditional learning algorithms with explicit training objectives, ICL enables LLMs to learn from demonstrations without explicit optimization (Brown et al., 2020; Min et al., 2022).
-Since traditional poisoning strategies are designed specifically to target the training process and exploit loss functions in conventional models (Biggio et al., 2012; Steinhardt et al., 2017; Geiping et al., 2020; He et al., 2023), the discrepancy between ICL and these strategies implies that they are not directly applicable to ICL.
-Conducting effective data poisoning attacks for ICL demands a thorough understanding of the unique learning mechanism of ICL.
+Unlike the traditional data poisoning attack (Steinhardt et al., 2017) where attackers manipulate the training data to corrupt the model trained on the poisoned data, we focus on the perturbations in examples utilized in ICL while keeping the LLMs untouched.
+This paper aims to answer the above question by exploring data poisoning attacks in ICL and uncovering the vulnerability of ICL when facing these attacks.
+We consider the standard pipeline of ICL, where examples are randomly selected from a data set for the corresponding downstream task.
+In terms of the attack, we assume that the adversary deliberately inserts some poisoned examples in this data, and their goal is to ensure that the learning process is adversely affected and the overall ICL prediction performance decreases.
+This scenario can be both significant and practical.
+For instance, LLMs and ICL are used in medical systems (Shi et al., 2024) for multiple tasks including diagnosis and cost computation.
+Healthcare providers such as hospitals and physicians may manipulate electronic health records (EHR) to increase revenue (Finlayson et al., 2018).
+Technically, data poisoning in ICL faces both unique challenges specific to ICL and common obstacles in traditional data poisoning.
+First, in contrast to traditional learning algorithms with explicit training objectives, ICL enables LLMs to learn from demonstrations without explicit training (Brown et al., 2020; Min et al., 2022).
+Since traditional poisoning strategies are designed specifically to target the training process and exploit loss functions in conventional models (Biggio et al., 2012; Steinhardt et al., 2017; Geiping et al., 2020; He et al., 2023), they are not directly applicable to ICL.
+Conducting effective data poisoning attacks for ICL requires a thorough understanding of the unique learning mechanism of ICL.
 Second, similar to traditional attacking methods, data poisoning for ICL also requires creating samples that are imperceptible to humans yet disruptive.
-These poisoned examples must seamlessly integrate with the other data to harm the learning process in a subtle way.
-While both traditional data poisoning and the attack in ICL require careful designs of the attack, for ICL, one extra challenge arises from the discrete vocabulary of language models, making it hard to manipulate inputs for effective disturbance (Lei et al., 2019; Xu et al., 2023).
-To tackle the above challenges, we introduce a novel and versatile attacking framework, ICLPoison, to exploit the unique learning mechanism of ICL.
-In particular, previous research (Xie et al., 2021; Hendel et al., 2023; Liu et al., 2023b; Wang et al., 2023) has shown a strong correlation between ICL performance and the hidden states within LLMs.
-Our framework, ICLPoison, ingeniously distorts these hidden states through strategic text perturbations—a non-trivial adaption given the attacker’s limited ability to only alter the examples in the demonstration.
-We further design three strategies for instantiating and optimizing poisoning attacks under the ICLPoison framework.
-Comprehensive experiments across various LLMs and tasks demonstrate the effectiveness of our methods, highlighting the vulnerability of ICL.
+These poisoned examples must seamlessly integrate with the other data to harm the learning process subtly.
+Moreover, one extra challenge of poisoning ICL arises from the discrete vocabulary of language models, making it hard to manipulate inputs for effective disturbance (Lei et al., 2019; Xu et al., 2023).
+To tackle the above challenges, we introduce a novel and versatile attacking method, ICLPoison, to exploit the unique learning mechanism of ICL.
+In particular, previous research (Xie et al., 2021; Hendel et al., 2023; Liu et al., 2023; Wang et al., 2023) has shown a strong correlation between ICL performance and the hidden states within LLMs.
+ICLPoison distorts these hidden states through strategic text perturbations, despite the attacker’s limited ability to alter only the demonstration examples.
+We further design three strategies for instantiating and optimizing poisoning attacks under the ICLPoison method.
+Extensive and comprehensive experiments across various LLMs and tasks demonstrate the effectiveness of our methods, highlighting the vulnerability of ICL.
 Notably, we have successfully degraded the performance of ICL in advanced models, including GPT-4 (a 10% decrease in ICL accuracy).
-Our study significantly advances the understanding of ICL’s vulnerabilities to data poisoning, bolstering the security and reliability of LLMs.
+Our study enhances the understanding of ICL’s vulnerability to data poisoning, which helps improve the security and reliability of LLMs.
+2 Related Works
+2.1 Data Poisoning Attack
+Data poisoning attacks (Biggio et al., 2012; Steinhardt et al., 2017) usually occur during the data collection phase of machine learning model training, where the training data is tampered with to induce malicious behaviors in the resulting models (He et al., 2023).
+These malicious behaviors include causing a poisoned model to have a poor overall accuracy (Steinhardt et al., 2017; Fowl et al., 2021; Huang et al., 2021), misclassifying a specified subset of test samples (Shafahi et al., 2018; Zhu et al., 2019), or inserting backdoors (Chen et al., 2017; Gu et al., 2019).
+The data poisoning on traditional NLP models is widely explored.
+CARA (Chan et al., 2020) inserts backdoors into text classifiers by generating poisoned samples by utilizing a conditional adversarially regularized autoencoder; (Yang et al., 2021; Schuster et al., 2020) poison the embeddings space of NLP models; (Chen et al., 2021) investigates different triggers to evaluate the effectiveness of backdoor attacks on NLP tasks.
+With the development of LLMs, poisoning attacks have also drawn the attention of the whole community, and studies have examined the vulnerability of LLMs against corruption in the pre-training and fine-tuning data.
+For example, BadPrompt (Cai et al., 2022) inserts backdoors during prompt-tuning and selects effective triggers to maximize the poisoning effect; POISONPROMPT (Yao et al., 2024) focuses on a similar setting and leverages a bi-level optimization objective to optimize the trigger while maintaining utility on clean samples; NOTABLE (Mei et al., 2023) considers the transferability of backdoors and injects backdoors into the encoder to adapt the attack to different downstream tasks and prompting strategies.
+(Wan et al., 2023) further explores the threat during instruction tuning of LLMs and shows that attackers can manipulate model predictions whenever a desired trigger phrase appears in the input.
+However, these attacks still focus on altering the pre-training or fine-tuning data, and the threat of poisoning examples in ICL remains unexplored.
+2.2 Attacking In-context Learning (ICL)
+ICL is considered an efficient way to adapt downstream tasks with a few examples, thus the potential safety issue with regard to ICL is of great importance.
+For instance, (Qiang, 2024) optimizes adversarial suffixes of examples to mislead the generation of LLMs; (Wei et al., 2023) also employs harmful demonstrations to induce LLMs to produce harmful responses.
+(Kandpal et al., 2023) first explores the threat of backdoor attacks on ICL.
+In particular, they insert backdoors in pre-training LLMs via fine-tuning on a poisoned dataset.
+Then when the prompt consists of examples from the target task, the backdoored LLM will misclassify the test sample containing the trigger as the malicious label.
+Following works (Zhao et al., 2024a, 2023) propose to directly add triggers into the demonstrations to avoid the computation cost of pre-training or fine-tuning.
+However, these works require the insertion of obvious triggers that can be easy to detect.
+In this work, we study more practical scenarios in which imperceptible perturbations are added to demonstrations, and attackers aim to compromise the ICL’s overall effectiveness.
+This presents unique challenges due to the implicit learning mechanism of ICL.
 3 Preliminary
-In this section, we introduce some background of ICL.
-We then present the preliminary results that the distortion in hidden states indeed affects the ICL performance, which inspires our framework.
-Key notations are also introduced.
-2.1 In-context Learning (ICL)
-[…]
-2.2 Understanding of ICL via Hidden States
-The emergence of ICL ability in LLMs has gained substantial attention, particularly in understanding its working mechanisms (Xie et al., 2021; Hendel et al., 2023; Von Oswald et al., 2023; Garg et al., 2022; Bai et al., 2023).
-Researchers have demonstrated that during ICL, LLMs can effectively extract “latent concepts” (denoted as $\theta$) from demonstrations (Xie et al., 2021; Wang et al., 2023), and the ICL predicting procedure can be decomposed as
-$P(y^{query}|[S,x^{query}])$ (1)
-[…]
-where $P(\cdot)$ denotes the probability distribution during the generation process of LLMs.
-In other words, ICL learns $\theta$ from the demonstration first and then make a prediction for $x^{query}$.
-The latent concept $\theta$ characterizes the intricate patterns and contextual information embedded in the data regarding the task.
-A key in this process is the hidden states, represented as $h$, which are defined as the representations of the last token of the input prompt at different layers of the model (as indicated by various studies (Hendel et al., 2023; Liu et al., 2023b; Todd et al., 2023)).
-The hidden states are used in encoding the latent concepts.
+In this section, we introduce ICL and its hidden states, along with the necessary notations.
+ICL.
+ICL is a paradigm that allows LLMs to learn tasks given a few examples in the form of demonstrations (Brown et al., 2020).
+To mathematically define the notation of ICL, suppose that we have a pre-trained LLM $f$, which takes in an input prompt $p$ and outputs a response $y$, i.e.
+$f(p)=y$ [^1] Since the main focus of this paper is not on the generation of LLMs, we adopt the default generation scheme for each LLM..
+Given a task $t\in\mathcal{T}$ from the ICL task set $\mathcal{T}$, we assume $(x,y)\sim\mathcal{D}_{t}$ where $x\in\mathcal{X}_{t}$ and $y\in\mathcal{Y}_{t}$.
+We further assume a prompt set $D_{t}=\{(x_{i,t},y_{i,t})\}_{i=1}^{N}$, where $(x_{i,t},y_{i,t})\sim\mathcal{D}_{t}$.
+To conduct an ICL prediction for a query $x^{query}\in\mathcal{X}_{t}$ under task $t$, the user first randomly selects $k$ input-output pairs from $D_{t}$ and concatenates them as a demonstration $S$, i.e.
+$S=[(x_{i,t},y_{i,t})]_{i=1}^{k}$.
+The demonstration is combined with query $x^{query}$ as an input prompt, and this prompt is sent to the LLM.
+We define the prediction of ICL as $\hat{y}^{query}_{ICL}=f([S,x^{query}])$.
+In this work, we consider classification tasks such as sentiment analysis and text classification.
+Hidden states of ICL.
+Extensive studies are conducted to understand the mechanisms of ICL (Xie et al., 2021; Hendel et al., 2023; Von Oswald et al., 2023; Garg et al., 2022; Bai et al., 2023).
+Researchers have demonstrated that the hidden states (represented as $h$), which are defined as the representations of the last token of the input prompt at different layers of the model (as indicated by various studies (Hendel et al., 2023; Liu et al., 2023; Todd et al., 2023)), plays a critical role in ICL.
+The hidden states can encode the latent concepts in the examples corresponding to the task, which further guides the prediction.
 In particular, consider a model $f$ composed of $L$ transformer layers, where each layer produces a vector representation $h_{l}(p,f)\in\mathbb{R}^{d}$ for the last token of the input prompt $p$ and $l\in[L]$.
-It has been observed that the hidden state $h_{l}$ at certain intermediate layers effectively works as $\theta$ (Hendel et al., 2023).
-Moreover, further research (Liu et al., 2023b; Todd et al., 2023) has been conducted to distill more concise representations from the set of hidden states $\{h_{l}\}_{l=1}^{L}$, enhancing the concept formation in ICL.
-2.3 Preliminary Experiments
-As discussed in the above, hidden states of the input prompt encode the latent concept and are closely related to the performance of ICL.
-Therefore, we conduct preliminary experiments to explore whether directly perturbing these hidden states can degrade ICL, and whether ICL is more vulnerable when perturbing the hidden states in all layers.
-[…]
-Figure 1: Results when directly perturbing the hidden states with Gaussian noises of various standard deviations (noise strength).
-3 A Novel Framework: ICLPoison
-In this section, we introduce a novel framework, ICLPoison, to conduct data poisoning by distorting hidden states of LLMs during the ICL process.
+It has been observed that the LLMs can make correct predictions conditioning on hidden states $h_{l}$ extracted from certain intermediate layers (Hendel et al., 2023).
+The hidden states of LLMs provide numerical representations and condense the information from the examples (Liu et al., 2023; Todd et al., 2023), which inspires our design of ICLPoison, as illustrated in the next section.
+4 ICLPoison
+In this section, we introduce a novel attacking method, ICLPoison, to conduct data poisoning by distorting hidden states of LLMs during ICL.
 4.1 Threat Model
-We assume that attackers have the full access to a portion of the dataset, specifically a subset related to a particular task.
-However, they may lack complete knowledge of the entire prompting set.
-Crucially, they are unaware of the details of the ICL process, including the test data, the Large Language Models (LLMs) employed, and specific ICL configurations like the number of examples and the templates used for demonstrations.
-Despite these limitations, the attackers can leverage open-source LLMs to introduce poisoned data into the subset they control.
+We assume the attacker’s goal is to compromise the ICL performance when adapting the downstream tasks using examples from the poisoned dataset as the demonstrations.
+We assume the attacker can insert poisoned data into the dataset.
+Crucially, attackers are unaware of the details of the ICL process, including the test data, the Large Language Models (LLMs) employed, and specific ICL configurations such as the number of examples and the templates used for demonstrations.
+Despite these limitations, the attackers can leverage open-source LLMs to generate poisoned data.
+In real practice, there are many possible scenarios where the attack can insert the poisoned data into the target data set.
+For instance, attackers can insert misinformation into the database used in a system where data is collected from various sources (Zou et al., 2024); the attackers can also manipulate third-party APIs to insert poisoned examples into the demonstrations (Zhao et al., 2024b).
+These threats bring safety and ethical concerns for the real applications of ICL in safety-critic domains like healthcare (Shi et al., 2024; Joe et al., 2021), finance (Loukas et al., 2023; Paladini et al., 2023), education (Li et al., 2024).
 4.2 ICLPoison Design
-As highlighted in Section 2.1, ICL differs from traditional learning algorithms since it lacks an explicit training objective that can be directly targeted by data poisoning attacks.
-To address this unique challenge, we draw inspiration from the dynamics of hidden states from Section 2.2, which indicates that the representation of the last token of the demonstration encodes the latent concept about the task.
-Building on this understanding, we introduce ICLPoison, a novel data poisoning framework specifically designed for the ICL process.
-ICLPoison strategically alters examples in demonstrations to distort the hidden states in the LLM, leveraging these internal dynamics to achieve our poisoning goals.
+As highlighted in Section 3, ICL differs from traditional learning algorithms since it lacks an explicit training objective that can be directly targeted by data poisoning attacks.
+To address this unique challenge, we draw inspiration from the understanding of hidden states in Section 3 and introduce ICLPoison, a novel data poisoning attack specifically designed for the ICL process.
+ICLPoison strategically alters examples in demonstrations to distort the hidden states for the poisoning goal.
 The details of ICLPoison are as follows.
-We focus on a surrogate LLM $f$ with $L$ layers and a task-specific prompt set $D^{p}_{t}=\{(x_{i,t}^{p},y_{i,t}^{p})\}_{i=1}^{N}$ from task $t$.
-Our objective is to diminish the ICL accuracy for the task $t$ by maximizing the distortion of hidden states in samples from $D^{p}_{t}$.
-We propose perturbing the input $x^{p}_{i,t}\sim\mathcal{X}_{t}$ using a transformation $\delta:\mathcal{X}_{t}\rightarrow\mathcal{X}_{t}$ while keeping the label $y^{p}_{i,t}$ unchanged.
+We focus on a surrogate LLM $f$ with $L$ layers and $(x,y)\sim\mathcal{D}_{t}$.
+Our objective is to reduce the ICL accuracy for the task $t$ by maximizing the distortion of hidden states of $(x,y)$.
+We propose perturbing the input $x\sim\mathcal{X}_{t}$ using a transformation $\delta:\mathcal{X}_{t}\rightarrow\mathcal{X}_{t}$ while keeping the label $y$ unchanged.
 The perturbation $\delta$ must be imperceptible to humans, hence it is constrained within a set $\Delta$ of imperceptible mappings.
 Details of $\delta$ will be discussed in Section 4.3.
-For each example $(x_{i,t}^{p},y_{i,t}^{p})$ in $D^{p}_{t}$, we extract its hidden states.
-Since the attacker lacks the knowledge of the test data, we use a dummy query $x^{query}_{t}\sim\mathcal{X}_{t}$ as a stand-in (as mentioned in (Hendel et al., 2023)).
-We then concatenate $(x_{i,t}^{p},y_{i,t}^{p})$ with $x^{query}_{t}$ to create a demonstration, and denote $h_{l}(x_{i,t}^{p},f)$ as the representation of the last token in the $l^{th}$ layer of model $f$.
-Since our focus is on perturbing $x^{p}_{i,t}$, we omit $y^{p}_{i,t}$ in the following discussion.
-The representations from all $L$ layers of model $f$ regarding input $x_{i,t}^{p}$ are denoted as $H(x_{i,t}^{p},f):=\{h_{l}(x_{i,t}^{p},f)\}_{l=1}^{L}$, representing the hidden states for $x^{p}_{i,t}$ under model $f$.
-For the perturbed input, the hidden states are $H(\delta_{i}(x_{i,t}^{p}),f)$, with $\delta_{i}$ being the specific perturbation for $x_{i,t}^{p}$.
-Our observations in Figure 1 indicate that altering the hidden state of just one layer is not optimal to significantly impact ICL performance.
-Therefore, we aim to maximize the minimum difference across all layers between the original and perturbed hidden states.
-To normalize differences across layers with varying scales, we use the normalized $L_{2}$ norm to measure the distance of the hidden state between the original example and the perturbed one for each layer: $l_{d}(h_{l}(x_{i,t}^{p},f),h_{l}(\delta_{i}(x_{i,t}^{p}),f))=\|\frac{h_{l}(x_{i,t}^{p},f)}{\|h_{l}(x_{i,t}^{p},f)\|_{2}}-\frac{h_{l}(\delta_{i}(x_{i,t}^{p}),f)}{\|h_{l}(\delta_{i}(x_{i,t}^{p}),f)\|_{2}}\|_{2}$.
-The distortion between $x_{i,t}^{p}$ and $\delta_{i}(x_{i,t}^{p})$ is further defined as:
-$\mathcal{L}_{d}(H(x_{i,t}^{p},f),H(\delta_{i}(x_{i,t}^{p}),f))$ (2)
-$=\min\limits_{l\in[L]}l_{d}(h_{l}(x_{i,t}^{p},f),h_{l}(\delta_{i}(x_{i,t}^{p}),f)).$
-The attacking objective of ICLPoison then becomes
-$\max\limits_{\delta_{i}\in\Delta}\mathcal{L}_{d}(H(x_{i,t}^{p},f),H(\delta_{i}(x_{i,t}^{p}),f)).$ (3)
-In other words, $\mathcal{L}_{d}(H(x_{i,t}^{p},f),H(\delta_{i}(x_{i,t}^{p}),f))$ denotes the minimum changes (or lower bound of the distortion) caused by the perturbation $\delta_{i}$ across all the layers of the model.
-This approach ensures that the perturbation $\delta_{i}$ introduces the most substantial change to the hidden states in the LLM during the ICL process, potentially enhancing the effectiveness of the poisoning attack.
+For each example $(x,y)$ possessed by the attacker, we extract its hidden states.
+Since the attacker lacks the knowledge of the test data, we use a dummy query $x^{query}_{t}\sim\mathcal{X}_{t}$ as a stand-in as suggested in (Hendel et al., 2023).
+We then concatenate $(x,y)$ with $x^{query}_{t}$ to create a demonstration and denote $h_{l}(x,f)$ as the representation of the last token in the $l^{th}$ layer of model $f$.
+Since our focus is on perturbing $x$, we omit $y$ in the following discussion.
+The representations from all $L$ layers of model $f$ regarding input $x$ are denoted as $H(x,f):=\{h_{l}(x,f)\}_{l=1}^{L}$, representing the hidden states for $x$ under model $f$.
+For the perturbed input, the hidden states are $H(\delta_{x}(x),f)$, with $\delta_{x}$ being the specific perturbation for $x$.
+To maximize the poisoning effect, we aim to maximize the minimum difference across all layers between the original and perturbed hidden states.
+To normalize differences across layers with varying scales, we use the normalized $L_{2}$ norm to measure the distance of the hidden state between the original example and the perturbed one for each layer: $l_{d}(h_{l}(x,f),h_{l}(\delta_{x}(x),f))=\|\frac{h_{l}(x,f)}{\|h_{l}(x,f)\|_{2}}-\frac{h_{l}(\delta_{x}(x),f)}{\|h_{l}(\delta_{x}(x),f)\|_{2}}\|_{2}$.
+The distortion between $x$ and $\delta_{i}(x)$ is further defined as:
+$\mathcal{L}_{d}(H(x,f),H(\delta_{x}(x),f))$ (1)
+$=\min\limits_{l\in[L]}l_{d}(h_{l}(x,f),h_{l}(\delta_{x}(x),f)).$
+The attacking objective of ICLPoison is
+$\max\limits_{\delta_{x}\in\Delta}\mathcal{L}_{d}(H(x,f),H(\delta_{x}(x),f)).$ (2)
+In other words, $\mathcal{L}_{d}(H(x,f),H(\delta_{x}(x),f))$ denotes the minimum changes (or lower bound of the distortion) caused by the perturbation $\delta_{x}$ across all the layers of the model.
+This approach ensures that the perturbation $\delta_{x}$ introduces the most substantial change to the hidden states in the LLM during ICL.
+By optimizing the objective in Equation 2 for each accessible example $(x,y)$, the attacker can create a poisoned example set $D^{p}_{t}$.
 4.3 Attacking Algorithms
-The perturbation $\delta$ is a crucial component of the ICLPoison framework, designed to be imperceptible to humans while effective in manipulating the performance.
-These are also common requirements for NLP attacks (Ebrahimi et al., 2017; Jin et al., 2020; Xu et al., 2023; Li et al., 2018).
-On the other hand, the discrete nature of the objective in Eq.3 poses additional challenges in the optimization.
-To address these challenges as well as showcase the versatility of our framework, we introduce three representative perturbations: synonym replacement, character replacement, and adversarial suffix.
-These methods demonstrate the adaptability of ICLPoison across different levels of text manipulation: Synonym replacement evaluates the word-level vulnerability of ICL and subtly changes the semantics; character replacement involves minimal but precise alterations, such as changing individual letters, making it less noticeable to human reviewers (see examples in Appendix A.3); and adversarial suffix test token-level vulnerabilities in ICL, accommodating different models by adapting to their unique tokenization methods.
-In general, these perturbations allow for a comprehensive evaluation of the vulnerability in ICL.
-The optimization of these perturbations is efficiently managed through a greedy search method, proven effective in similar contexts (Lei et al., 2019; Bao et al., 2022).
+To design the perturbation $\delta$, as required in common NLP attacks (Ebrahimi et al., 2017; Jin et al., 2020; Xu et al., 2023; Li et al., 2018), $\delta$ is supposed to be imperceptible to humans while effective in manipulating the performance.
+In addition to the above requirements, an additional challenge to consider in the optimization is the discrete nature of the objective in Eq.2.
+To address these requirements and challenges as well as showcase the versatility of our method, we introduce three representative perturbations: synonym replacement, character replacement, and adversarial suffix.
+These methods demonstrate the adaptability of ICLPoison across different levels of text manipulation: Synonym replacement evaluates the word-level vulnerability of ICL and subtly changes the semantics; character replacement involves minimal but precise alterations, making it less noticeable to human reviewers (see examples in Appendix A.4); and adversarial suffix test token-level vulnerabilities in ICL.
+The optimization of these perturbations is managed through a greedy search method, proven effective in similar contexts (Lei et al., 2019; Bao et al., 2022).
 Synonym Replacement.
 This approach involves substituting words in a text with their synonyms, aiming at preserving the semantic meaning and grammatical structure (Jin et al., 2020; Xu et al., 2023).
-Therefore, it serves as an appropriate example of $\delta$ in Section 3.2.
-Within our framework, we limit the number of word replacements (denoted as $k$) to maintain the perturbation’s imperceptibility.
+Within our method, we limit the number of word replacements (denoted as $k$) to maintain the perturbation’s imperceptibility.
 For a text composed of a sequence of $n$ words $x=[w_{1},...,w_{n}]$, $\delta(x)$ is defined as $[s_{1},...,s_{n}]$, where $s_{i}$ is either a synonym of $w_{i}$ (if selected for replacement) or remains as $w_{i}$ (if not replaced).
-Our objective is to identify the optimal replacements that maximize the objective in Eq.3.
-This brings natural challenges: which words are to be replaced and what synonyms are suitable.
-Following a strategy similar to (Jin et al., 2020), we adopt a two-step optimization process.
-Firstly, we calculate an importance score for each word, selecting the top-$k$ words with the highest scores for replacement.
+To identify which words are to be replaced and which synonyms are suitable, we follow a strategy similar to (Jin et al., 2020).
+We adopt a two-step optimization process.
+First, we calculate an importance score for each word, selecting the top-$k$ words with the highest scores for replacement.
 The importance score for word $w_{i}$ is computed as the distortion (in Eq.1) before and after deleting $w_{i}$, expressed as:
@@
 where $x_{\text{\textbackslash}w_{i}}$ denotes the text after the removal of $w_{i}$.
-In the second step, we use a greedy search method.
-This method involves iteratively finding the best replacement for each selected word, one at a time, while keeping the other words fixed (Yang et al., 2020; Lei et al., 2019).
-In each iteration, we focus on one of the selected words, replacing it with its most appropriate synonym.
-We retrieve synonyms using GloVe word embeddings (Pennington et al., 2014), selecting those with embeddings most similar (based on cosine similarity) to the original word.
-After identifying potential synonyms, each synonym is temporarily substituted into the text, and the loss function in Eq.3 is evaluated.
-The synonym that results in the highest loss is then chosen as the final replacement.
-We then proceed to the next word, repeating this replacement process until traversing all selected words.
+In the second step, we use a greedy search method that iteratively replaces each selected word while keeping others fixed (Yang et al., 2020; Lei et al., 2019).
+For each word, we find synonyms using GloVe embeddings (Pennington et al., 2014), selecting those with the highest cosine similarity to the original word.
+Each synonym is temporarily substituted into the text, and the loss function in Eq.2 is evaluated.
+The synonym that maximizes the loss is chosen as the final replacement.
+This process is repeated for all selected words.
 We present the detailed algorithm in Algorithm.
@@
 Character Replacement.
-This method is similar with the synonym replacement approach, but focuses on replacing individual characters instead of whole words (Ebrahimi et al., 2017; Lei et al., 2019; Xu et al., 2023).
+This method is similar to the synonym replacement approach but focuses on replacing individual characters instead of whole words (Ebrahimi et al., 2017; Xu et al., 2023).
 When changing only a few letters, this method can be less detectable to humans and maintain the word’s pronunciation and basic structures (Ebrahimi et al., 2017).
-We limit the number of character replacements to $k$ to maintain the subtlety of the perturbations.
-The optimization process for character replacement also follows a two-step method.
-Firstly, we calculate the importance score for each character, similar to the approach described in Eq.4.
-The primary difference is that we consider the removal of individual characters, not words.
-The top-$k$ characters with the highest importance scores are earmarked for replacement.
-Secondly, we employ the same greedy search strategy used in synonym replacement to identify the optimal replacements for the selected characters.
+We limit character replacements to $k$ to keep perturbations subtle.
+The optimization process involves two steps: first, we calculate an importance score for each character, similar to Eq.3 but focusing on individual character removal rather than words.
+The top-$k$ characters with the highest scores are selected for replacement.
+Second, we use a greedy search strategy, similar to synonym replacement, to find the best replacements for these characters.
 Note that our character set encompasses uppercase and lowercase letters, digits, punctuation marks, and whitespace, in line with the sets in (Kim et al., 2016; Ebrahimi et al., 2017).
@@
 Thus, in addition to synonym and character replacement, we also adapt this perturbation to evaluate the token-level vulnerability of the ICL process.
-To ensure imperceptible to humans, we restrict the number of additional tokens when adapting to our framework.
+To ensure imperceptible to humans, we restrict the number of additional tokens when adapting to our method.
 For a given text $x$ that can be tokenized into a sequence of tokens $x=[t_{1},...,t_{n}]$, we define $\delta(x)$ as $[t_{1},...,t_{n},t^{\prime}_{1},...,t^{\prime}_{k}]$ where $t^{\prime}_{1},...,t^{\prime}_{k}$ are adversarial suffices.
 Our goal is to identify the optimal suffixes that maximize the objective in Eq 2.
-We also employ a greedy search approach, involving iteratively selecting each suffix token from $t^{\prime}_{1}$ to $t^{\prime}_{k}$ one by one which results in the maximum increase in the loss.
-The detailed implementation and optimization process for this approach is elaborated in Algorithm 3 in Appendix  A.1.
-4 Experiment
-We conduct extensive experiments to validate the effectiveness of the proposed framework ICLPoison, particularly with three perturbations introduced in Section 3.3.
+We also employ a greedy search approach and iteratively select each suffix token from $t^{\prime}_{1}$ to $t^{\prime}_{k}$ one by one which results in the maximum increase in the loss.
+The detailed implementation and optimization process for this approach is in Algorithm 3 in Appendix A.1.
+5 Experiments
+We conduct extensive experiments to validate the effectiveness of the proposed method ICLPoison, particularly with three perturbations introduced in Section 4.3.
 5.1 Experiments setting
@@
 We conduct experiments on different types of classification tasks.
-Stanford Sentiment Treebank (SST2) dataset from the GLUE benchmark (Wang et al., 2019) is a sentiment analysis dataset consisting of sentences from movie reviews and human annotations of their sentiment in 2 classes (positive/negative); Corpus of Linguistic Acceptability (Cola) dataset from GLUE is a linguistic analysis dataset consisting of English acceptability judgments collected from linguistic books, labeled with “acceptable” or “unacceptable”; Emo dataset (Wang et al., 2023) focuses on emotion classification consisting of Twitter messages labeled in 4 classes; AG’s new (AG) corpus (Zhang et al., 2015) is a topic classification dataset gathered from news sources and labeled in 4 classes (World, Sports, Business, and Science/Technology); and Poem Sentiment (Poem) (Sheng & Uthus, 2020) is a sentiment analysis dataset of poem verses from Project Gutenberg, classified into 3 classes.
+SST2 (2 classes) and Cola (2 classes) are from GLUE dataset (Wang et al., 2019); Emo (Wang et al., 2023) (4 classes) is an emotion classification dataset; AG’s new (4 classes) (Zhang et al., 2015) is a topic classification dataset; Poem (3 classes) (Sheng and Uthus, 2020) is a sentiment analysis dataset of poem.
+Details of these datasets are presented in Appendix A.2.
 Models.
-In this work, we consider both open-source models including Llama2-7B (Touvron et al., 2023), Pythia (2.8B, 6.9B) (Biderman et al., 2023), Falcon-7B (Almazrouei et al., 2023), GPT-J-6B (Wang & Komatsuzaki, 2021), MPT-7B (Team, 2023), and API-only models GPT-3.5 and GPT-4 (Brown et al., 2020).
+We use open-source models including Llama2-7B (Touvron et al., 2023), Pythia (2.8B, 6.9B) (Biderman et al., 2023), Falcon-7B (Almazrouei et al., 2023) [^2] Two additional models GPT-J-6B (Wang and Komatsuzaki, 2021), MPT-7B (Team, 2023) in Appendix A.3, and API-only models GPT-3.5 and GPT-4.
 Baselines.
@@
 Metrics.
-Our main focus is on the ICL accuracy.
+Our main focus is on ICL accuracy.
 For every dataset, we generate perturbations for examples in the training data and randomly select examples from it to conduct ICL prediction for every sample in the test data.
@@
 Falcon-7B AG 75.2$\pm$1.8 50.8$\pm$1.3 11.2$\pm$2.3 14.9$\pm$1.7 12.8$\pm$1.2
-Cola 57.8$\pm$1.3 49.1$\pm$2.5 13.7$\pm$1.7 17.2$\pm$1.8 11.8$\pm$0.9
-SST2 85.4$\pm$1.6 82.8$\pm$2.1 14.8$\pm$2.0 18.9$\pm$1.5 11.4$\pm$1.1
-Emo 58.7$\pm$1.1 46.2$\pm$1.7 11.7$\pm$1.8 13.8$\pm$1.3 9.6$\pm$0.7
-Poem 57.6$\pm$1.5 46.7$\pm$1.4 12.6$\pm$2.4 14.2$\pm$2.2 10.3$\pm$1.3
-GPT-J-6B AG 63.2$\pm$1.7 53.4$\pm$1.9 11.9$\pm$1.5 16.8$\pm$1.8 12.5$\pm$1.1
 Cola 75.6$\pm$0.7 76.3$\pm$0.6 58.1$\pm$0.5 62.6$\pm$0.4 59.7$\pm$0.4
@@
 GPT-4 AG 90.4$\pm$0.3 87.3$\pm$0.3 83.2$\pm$0.5 83.1$\pm$0.4 84.7$\pm$0.5
-4.2 Main results
-In this subsection, we apply our framework to various LLMs and datasets to evaluate the vulnerability of ICL.
-Attacking open-source models.
-Our study initially examines open-source models.
-For each, we craft poisoned samples utilizing the model’s own architecture and subsequently assess the ICL accuracy.
-Note that we will discuss the transferability i.e.
-examining the performance of poisoned samples crafted using Llama2-7B when tested on different models in Section 4.3.
-We begin by altering the entire training set to create these samples.
-Further studies involving only a subset of poisoned data are presented in Section 4.4.
-Partial results are shown in Table 1 and full results can be found in Table 5 in Appendix A.2.
-In the table, a lower accuracy indicates a stronger poisoning effect and the lowest performance is highlighted.
-It is obvious that ICL performs well on clean data, especially for the SST2 dataset and Llama2-7B model, achieving more than 88% accuracy.
-While random labels do interrupt the ICL process, causing a decrease in performance, LLMs are still capable of learning effectively from the demonstrations with the dropping of ICL accuracy less than 7%, which is aligned with observations by (Min et al., 2022).
-In contrast, our ICLPoison framework significantly reduces ICL accuracy, achieving drops to below 10% for some models and datasets such as GPT-J-6B with the Emo dataset.
-Notably, ICLPoison reveals that ICL is vulnerable to different levels of data poisoning attacks, ranging from subtle character-level changes to word-level manipulations.
-Among all three variants of ICLPoison, synonym replacement and adversarial suffix perturbations exhibit more severe threats than character replacement in reducing ICL accuracy, likely because they introduce more changes within the same perturbation budget.
-We present some poisoned examples from three methods in Appendix A.3 for human evaluation.
-We observe that the poisoning effect differs across models and datasets.
-For instance, Llama2-7B shows heightened sensitivity to poisoning on the Emo dataset, whereas Falcon-7B is particularly vulnerable to attacks on the Poem Sentiment dataset.
-These variations may stem from differences in model architectures and the inherent complexity of the datasets.
+5.2 Effectiveness of ICLPoison
 Figure 1: Experimental results of transferring poisoned data from Llama2-7B to other models.
 The Y-axis represents the ICL accuracy (a smaller value represents a stronger poisoning effect), while the X-axis denotes different models.
+In this subsection, we first validate the effectiveness of the idea of corrupting the hidden states in ICLPoison across various LLMs and tasks.
+To eliminate the potential impact on random sampling of poisoned/clean examples, we consider the extreme case of 100% poisoning rate [^4] This is an extreme case but still practical.
+For instance, the victim downloads a whole poisoned data set generated by the attacker for ICL prediction..
+A more practical attacking scenario with small poisoning rates is discussed in Section 5.3.
+Attacking open-source models.
+We first examine open-source models.
+For each, we craft poisoned samples utilizing the model’s own architecture and assess the ICL accuracy.
+Partial results are shown in Table 1 and full results can be found in Table 7 in Appendix A.3.
+In the table, a lower accuracy indicates a stronger poisoning effect, and the lowest performance is highlighted.
+One can see that ICL performs well on clean data, especially for the SST2 dataset and Llama2-7B model, achieving more than 88% accuracy.
+Besides, the random labeling can only decrease the accuracy marginally, less than 7%, which is aligned with observations by (Min et al., 2022).
+In contrast, our ICLPoison method significantly reduces ICL accuracy, achieving drops to below 10% for some models and datasets such as Falcon-7B with the Emo dataset.
+Notably, the effectiveness of ICLPoison implies that ICL is vulnerable to data poisoning attacks that corrupt the hidden states.
+In addition, among the three variants of ICLPoison, synonym replacement and adversarial suffix perturbations pose more severe threats to ICL accuracy compared to character replacement.
+This disparity may arise because, within the same perturbation budget, different types of perturbations induce varying degrees of change in the hidden states.
+Character changes typically do not alter the semantic content as significantly as synonym replacements or adversarial suffixes, which can introduce more substantial shifts in the hidden states and disrupt the model’s prediction more effectively.
+We present some poisoned examples from three methods in Appendix A.4 for human evaluation.
 Attacking API-only models.
@@
 Notably, GPT-4 exhibits particular susceptibility to character replacement, suggesting a heightened sensitivity to minor textual variations.
-4.3 Transferbility
-In this subsection, we consider a practical scenario when the attacker has no access to the victim LLM, and leverages a surrogate model to generate poisoned samples.
-Specifically, we adopt the Llama2-7B model in this study.
-This setup allows us to evaluate the transferability of our poisoning approach across different models, including black-box models such as GPT-3.5 and GPT-4.
-The initial results for the transferability study are presented in Figure 2, in which we focus on the SST2 dataset.
-For a more comprehensive analysis covering all five datasets used in our study, we direct readers to Appendix A.2.
-From Figure 2, we observe a notable trend: while the efficacy of the poisoning attack diminishes when moving from the surrogate (Llama2-7B) to the victim models, the impact remains significant.
-The poisoned examples generated by our ICLPoison framework – in all three of its variants – substantially jeopardize the performance of ICL, leading to over a 30% decrease in accuracy for open-source models.
-Furthermore, our analysis sheds light on the different transferability of various perturbations.
-Synonym replacement and adversarial suffixes, in particular, demonstrate a stronger poisoning effect compared to character replacement.
-This disparity could be attributed to the more pronounced influence these methods exert on the surrogate model.
-We also note the influence of model size on susceptibility to poisoning.
-Models smaller in size than the surrogate, such as Pythia-2.8B and GPT-J-6B, appear to be more vulnerable to these poisoning examples, while larger models exhibit a degree of resistance.
-This pattern suggests that the effectiveness of our approach may be modulated by the size and complexity of the target model, offering valuable insights for future research in trustworthy language models.
-4.4 Partial poisoning
-To consider practical scenarios where attackers can only alter a part of the data, our experiments involve perturbing a random subset of the training dataset.
-We allocate varying proportions (10%, 20%, 50%, and 100%) of the training data as accessible for manipulation in the Llama2-7B model and GLUE-SST2 dataset.
-The results, shown in Figure 3, reveal that a lower mixing rate reduces the overall poisoning effect.
-Nonetheless, even at a 10% rate, we observe a significant decrease in ICL performance by over 7%, and a 15% decrease at 20%, underscoring the efficacy of our ICLPoison framework.
-Figure 3: Results for partially poisoned ICL. The X-axis represents the ratio of poisoned data while Y-axis represents the ICL accuracy.
+Transferbility.
+We adopt the Llama2-7B model as the surrogate model and test ICL performance on other models to evaluate the transferability of our poisoning approach across different models, including black-box models.
+This evaluation validates the transferability of proposed ICLPoison and reveals factors influencing the transferability such as perturbation type and model size.
+Initial results, presented in Figure 1, focus on the SST2 dataset, with a comprehensive analysis of all five datasets available in Appendix A.3.
+The results indicate that while the effectiveness of the poisoning attack diminishes when moving from the surrogate (Llama2-7B) to other models, the impact remains significant.
+Our ICLPoison method—across its three variants—leads to over a 30% decrease in accuracy for open-source models.
+This underscores the substantial threat posed by these attacks.
+Our analysis also reveals differences in how various perturbations transfer across models.
+Synonym replacement and adversarial suffixes demonstrate a stronger poisoning effect compared to character replacement, likely because they introduce more disruption to the model’s hidden states.
+Furthermore, smaller models, such as Pythia-2.8B and GPT-J-6B, are more susceptible to these poisoning examples, whereas larger models exhibit some resistance.
+This suggests that the effectiveness of our approach can be affected by the size and complexity of the target model.
+Figure 2: Results for practical ICLPoison. The X-axis represents the poison rate while Y-axis represents the ICL accuracy.
+The above experimental results confirm that distortions in hidden states can significantly compromise ICL predictions.
+Synonym replacements and adversarial suffixes cause greater changes in hidden states, leading to stronger poisoning effects.
+These distortions are also transferable, with smaller models being more affected.
+These insights underscore the critical link between hidden states and ICL predictions, emphasizing the effectiveness of data poisoning via manipulating hidden states.
+5.3 Practical Attacks
+To align with the threat model in Section 4.1, we consider practical attacking scenarios when attackers only insert some poisoned data into datasets.
+This usually happens when the victim collects data from various sources and mixes the poisoned data uploaded by the attacker with other clean data (Steinhardt et al., 2017; Carlini et al., 2024).
+In our experiments, we mix the poisoned data with clean training data, and the poisoning rate varies (10%, 20%, 50%, and 100%).
+We test with the Llama2-7B model and GLUE-SST2 dataset.
+The results, shown in Figure 2, reveal that ICLPoison is still a threat even when only a small set of poisoned data is injected into the clean prompt set.
+At a small poisoning rate such as 10% and 20%, we can observe a significant decrease in ICL performance by over 10% and 19% respectively, underscoring the efficacy of our proposed ICLPoison.
 These findings highlight the vulnerability of ICL to subtle data poisoning attacks, where even a limited number of malicious inputs can significantly disrupt the ICL process.
-Additionally, the results show that perturbation strategies like synonym replacement and adversarial suffix, which introduce more pronounced textual changes within a fixed poisoning budget, more severely affect ICL performance compared to character replacement.
-This indicates that broader, coarse-grained perturbations are generally more disruptive than finer, more subtle ones.
+Additionally, similar to the observations in Table 1, the results show that perturbation strategies like synonym replacement and adversarial suffix, which introduce more pronounced textual changes within a fixed poisoning budget, more severely affect ICL performance compared to character replacement.
+5.4 Potential Defenses
+Since the attackers are not aware of how victims leverage the poisoned data and some naive defenses may be applied, it is necessary to evaluate the robustness of an attack method.
+To examine the robustness of ICLPoison, we evaluate three representative defenses: perplexity filtering, paraphrasing (Jain et al., 2023), and grammar check.
+We also take 100% poisoning rate as the illustration and apply each defense.
 Table 2: Average perplexity scores for clean and poisoned data with standard error reported. Focus on model Llama2-7B.
@@
 AG 2.37$\pm$0.41 3.12$\pm$0.29 3.84$\pm$0.50 3.68$\pm$0.35
-4.5 Potential defenses
-To evaluate the robustness of our framework, we applied two representative defenses from (Jain et al., 2023): detection-based defense perplexity filter and preprocessing defense paraphrasing.
-Perplexity filter.
-The perplexity score of a text is referred to as the average negative log-likelihood of each of the tokens appearing.
-The perturbation added to the original text may cause grammar mistakes, logic problems, and a reduction in fluency.
-This will increase the perplexity, and thus the corresponding text could be more detectable (Jain et al., 2023).
-We report the perplexity scores of generated poisoned data, and a higher value indicates a higher probability of being detected and lower robustness.
-Table 2 shows the perplexity scores for poisoned data across various models and datasets, calculated as described in Section 4.1 of (Jain et al., 2023).
-Among the perturbations, synonym replacement yields relatively lower perplexity most close to the clean perplexity, whereas adversarial suffixes generally result in the highest scores.
-This suggests that synonym replacement is less detectable than the adversarial suffix.
+Perplexity filtering.
+The perplexity score measures the average negative log-likelihood of tokens in a text.
+Perturbations to the original text can increase this score, making them potentially detectable (Jain et al., 2023).
+We present the perplexity scores of generated poisoned data using the Llama2-7B model; higher scores suggest a greater chance of detection and defense.
+Table 2 displays these scores for poisoned data across various models and datasets, following the calculation method in Section 4.1 of (Jain et al., 2023).
+Among perturbations, synonym replacement results in perplexity scores closest to clean data, while adversarial suffixes yield the highest scores.
 Paraphrasing.
-Paraphrasing is a typical preprocessing method that utilizes a language model to rewrite the input text.
-This process aims to maintain original meanings while removing adversarial perturbations, making it an effective defensive strategy.
-In this part, we leverage the cutting-edge GPT-4 model to paraphrase the poisoned data and evaluate ICL performance on paraphrased inputs.
-As shown in Table 3, paraphrasing successfully neutralizes adversarial suffixes while largely preserving the impact of synonym replacements.
-This outcome is anticipated since adversarial suffixes typically introduce irrelevant content, whereas synonym replacements maintain the text’s semantic integrity.
-Table 3: The poisoned data is paraphrased by GPT-4, and ICL accuracy is reported. Original results are included in brackets.
+Paraphrasing is a preprocessing technique that uses a language model to rewrite text, aiming to preserve original meanings while removing adversarial perturbations.
+We use GPT-4 to paraphrase poisoned data and assess ICL performance on these inputs with the Llama2-7B model.
+Table 3 shows that paraphrasing effectively neutralizes adversarial suffixes but largely retains the effects of synonym replacements.
+To explain this, adversarial suffixes add irrelevant content, while synonym replacements preserve semantic integrity [^5] Paraphrased text examples are provided in Appendix A.5..
+Table 3: ICL accuracy when paraphrased by GPT-4. Original results are included in brackets.
 Datasets Clean Random label Synonym Character Adv suffix
@@
 AG 70.6(73.2) 55.7(57) 38.2(13.6) 47.2(19.4) 64.3(11.9)
-The above findings make it evident that the implemented defenses can enhance the robustness of ICL against data poisoning to some extent.
-However, the degree of improvement is contingent upon the type of perturbation.
-For instance, token-level perturbations, such as adversarial suffixes, are more readily addressed by these defenses, whereas word-level perturbations like synonym replacements pose a greater threat due to their subtlety and ability to alter semantic meaning.
-Character-level perturbations fall between these two in terms of detectability and potential impact, with their severity likely to increase with a higher attack budget.
-Our results highlight the necessity to develop a robust ICL.
+These findings indicate that defenses enhance ICL’s robustness against data poisoning, but their effectiveness varies by perturbation type.
+Token-level perturbations like adversarial suffixes are easier to counter, whereas word-level synonym replacements are more challenging due to subtle semantic changes.
+Character-level perturbations are moderately detectable but become more severe with larger attack budgets.
+These results highlight the need for more robust ICL defenses.
+Grammar check defenses.
+We evaluate the performance of ICLPoison against typo-corrector (Pruthi et al., 2019) and popular grammar-checking tools–Grammarly and Microsoft Editor.
+We conduct experiments on SST2 dataset and Llama2-7B model with all three attacking strategies.
+For typo-corrector, we report the average ICL accuracy; for Grammarly, we report the ratio of correctness errors (CER) over the total number of words; for Microsoft Editor, we report the ratio of spelling errors (SER) and the ratio of grammar errors (GER) over the total number of words.
+As shown in Table 4, the proposed attack can bypass the grammar/typo detection method, especially the synonym replacement strategy.
+This highlights the stealthiness of ICLPoison.
+Table 4: Attacking performance against detection defenses on model Llama2-7B and SST2 dataset.
+Defenses Clean Synonym Adv Suffix Character
+Grammarly
+(CER ($\downarrow$))
+4.89 5.04 11.97 10.31
+Microsoft editor
+(SER/GER ($\downarrow$))
+2.72/0.62 2.68/0.65 10.08/0.63 7.25/0.68
+Typo-corrector
+(ICL accuracy ($\downarrow$))
+90.2 23.7 41.4 35.8
 5.5 Ablation studies
-Under the assumption that attackers lack knowledge of the ICL process details, such as the templates and the number of examples used, we perform ablation studies to understand the impact of these settings.
-Specifically, we evaluate three different templates: F1, which is used in generating poisoned data; F2, formatted as ’input-output’; and F3, structured as ’Q:input, A:output’.
-Additionally, we explore the effects of using 3, 5, and 7 examples in ICL predictions.
-This ablation study concentrates on the Llama2-7B model and the GLUE-SST2 dataset, with comprehensive results provided in Appendix A.2.
-The results for different templates and varying numbers of examples are detailed in Table 4 and Figure 4, respectively.
-The results demonstrate that our framework is effectively adaptable to various templates and different numbers of demonstrations, highlighting its potentials in practice.
-Table 4: Experiments with different templates: F1, F2, F3 on model Llama2-7B and dataset GLUE-SST2.
-F1 F2 F3
-Clean 88.6 92.5 90.3
-Random label 82.1 84.7 79.2
-Synonym 18.5 17.9 18.2
-Character 26.8 30.6 28.5
-Adv suffix 20.4 21.7 19.3
-Figure 4: Experiments with different numbers of examples.
-5 Related Works
-5.1 In-context Learning
-In-context learning (ICL), since its introduction by (Brown et al., 2020), has been notable for efficiently tackling tasks with just a few examples, bypassing the need to adjust the model’s parameters (Dong et al., 2022; Liu et al., 2021; Sun et al., 2022).
-Generally speaking, ICL allows LLMs to make predictions based on a few examples formatted in a template rather than changing model parameters.
-Existing works have shown that the performance of ICL relies on the quality of demonstrations, including the selection of examples (Liu et al., 2021; Sorensen et al., 2022; Wang et al., 2023), the order of examples (Lu et al., 2021; Min et al., 2022), the template used to format demonstrations (Liu et al., 2023a; Wei et al., 2022).
-The mechanisms behind ICL’s success have been the subject of much interest.
-One line of work reveals that ICL learns latent concepts and conducts implicit Bayesian inference (Xie et al., 2021).
-Follow-up works (Hendel et al., 2023; Liu et al., 2023b) take a further step and point out that hidden states in LLMs encode these concepts.
-Another line of work focuses on the implicit learning mechanism of ICL.
-Garg et al.
-(2022) showed that Transformers (Vaswani et al., 2017) can encode effective learning algorithms to learn unseen linear functions according to demonstration samples.
-Von Oswald et al.
-(2023); Dai et al.
-(2022); Akyürek et al.
-(2022) connect Transformers with gradient descent and claim that LLMs can implicitly perform gradient descent on examples.
-Bai et al.
-(2023) further extend the ability of implicit gradient descent to various statistical abilities including pre-ICL testing and post-ICL validation.
-Our work builds on the foundation of the first line, and leverages the understanding of latent concepts and hidden states to conduct effective data poisoning attacks.
-5.2 Data Poisoning
-Data poisoning attacks (Biggio et al., 2012; Steinhardt et al., 2017) traditionally occur during the data collection phase of machine learning model training, where the training data is tampered with to induce malicious behaviors in the resulting models.
-These behaviors can range from degraded performance on testing data (Steinhardt et al., 2017; Huang et al., 2021) to the misclassification of specific test samples (Shafahi et al., 2018; Zhu et al., 2019) and the implantation of backdoors.
-In conventional models, such attacks typically exploit the training objectives (Steinhardt et al., 2017; He et al., 2023).
-However, applying data poisoning to in-context learning (ICL) aims to undermine the ICL’s overall effectiveness, presenting unique challenges due to its implicit learning mechanism.
+We present ablation studies with regard to the number of perturbations.
+Due to the page limit, we provide more ablation studies in Appendix A.3 including token length, prompt template, computation cost, etc.
+We vary the number of perturbations from 0 to 5, on the SST2 dataset and Llama2-7B model.
+As shown in Table 5, increasing the number of perturbations will enhance the poisoning effect.
+Moreover, even for a small number of perturbations, such as 2, we can significantly compromise the accuracy of ICL by at least 30%.
+Table 5: Ablation on the number of perturbations, model Llama2-7B and SST2 dataset.
+Budget 0(clean) 1 2 3 4 5
+Adversarial suffix 88.6 69.3 50.8 34.9 25.1 20.4
+Char replacement 88.6 73.5 57.2 46.8 32.7 26.8
+Synonym replacement 88.6 70.1 52.9 39.1 26.4 18.5
 6 Conclusion
-In this study, we introduce ICLPoison, a novel framework devised to assess the vulnerability of in-context learning (ICL) in the face of data poisoning attacks.
-We use the dynamics of hidden states in Large Language Models (LLMs) to craft our attack objectives.
-Furthermore, we implement our framework through three distinct and practical algorithms, each employing a different method of discrete perturbation.
+In this study, we introduce ICLPoison, a novel method devised to assess the vulnerability of ICL in the face of data poisoning attacks.
+We use the dynamics of hidden states in LLMs to craft our attack objectives.
+Furthermore, we implement our method through three distinct and practical algorithms, each employing a different method of discrete perturbation.
 Our research exposes previously unidentified susceptibilities of the ICL process to data poisoning.
 This discovery emphasizes the urgent need for enhancing the robustness of ICL implementations.
-Broader Impact
-This paper introduces a new framework to assess how in-context learning in large language models (LLMs) is susceptible to data poisoning attacks.
-It highlights the vulnerabilities in in-context learning, urging the development of more robust implementations.
-Additionally, the framework could aid in creating stronger attacks.
-Our findings underscore the pressing need to enhance LLMs’ security to protect against such vulnerabilities.
+7 Limitations
+In this work, we investigate the vulnerability of in-context learning (ICL) facing data poisoning attacks.
+In specific, we develop a method ICLPoison to evaluate the vulnerability.
+The proposed method needs access to the hidden states of LLMs, and thus can not be directly applied to API-only models.
+We only test on GPT-3.5 and GPT-4, and more API-only models need to be tested.
+Moreover, attacking methods that only require black-box access are to be investigated to directly attack API-only models.
+Acknowledgement
+Pengfei He and Jiliang Tang are supported by the National Science Foundation (NSF) under grant numbers CNS2321416, IIS2212032, IIS2212144, IOS2107215, DUE2234015, CNS2246050, DRL2405483 and IOS2035472, the Army Research Office (ARO) under grant number W911NF-21-1-0198, Amazon Faculty Award, JP Morgan Faculty Award, Meta, Microsoft and SNAP.
 Appendix A Appendix
@@
   Return poisoned prompting set $\{(\delta_{i}(x^{p}_{i,t}),y^{p}_{i,t})\}_{i=1}^{N}$
-A.2 Additional Experiments
+A.2 Details of datasets
+  • Stanford Sentiment Treebank (SST2) dataset from the GLUE benchmark (Wang et al., 2019) is a sentiment analysis dataset consisting of sentences from movie reviews and human annotations of their sentiment in 2 classes
+  • Corpus of Linguistic Acceptability (Cola) dataset from GLUE is a linguistic analysis dataset consisting of English acceptability judgments collected from linguistic books, labeled with “acceptable" or “unacceptable"
+  • Emo dataset (Wang et al., 2023) focuses on emotion classification consisting of Twitter messages labeled in 4 classes
+  • AG’s new (AG) corpus (Zhang et al., 2015) is a topic classification dataset gathered from news sources and labeled in 4 classes
+  • Poem Sentiment (Poem) (Sheng and Uthus, 2020) is a sentiment analysis dataset of poem verses from Project Gutenberg, classified into 3 classes
+A.3 Additional Experiments
 In this section, we present additional experimental results, including full results on attacking open-source models in Table 7, full results of transferability in Table 9, full results of perplexity scores in Table 11, full results on various templates in Table 8 and the number of examples in Table 10.
@@
 A detailed investigation can be an interesting future direction.
+Practical attacks. We provide the full results of practical attacks with different poison rates in Table 6.
+Table 6: Full results of practical attacks.
+Poisoning rate Clean Random label Synonym Character Adv suffix
+100% 88.6 82.1 18.5 26.8 20.4
+50% 88.6 84.6 46.3 50.1 43.5
+20% 88.6 87.5 67.1 70.8 65.9
+10% 88.6 88.1 75.6 77.5 73.2
 Perplexity scores. Table 11 covers perplexity scores on various datasets and models. It is obvious that synonym replacement is more stealthy than the other 2 methods.
@@
 GPT-4 AG 90.4 87.3 83.2 83.1 84.7
+Table 10: Full results for different number of examples in the demonstration. Focus on model Llama2-7B.
+Dataset num_examples Clean Random label Synonym Character Adv suffix
+Cola 3 63.2 59.2 16.5 21.8 12.8
+5 63.8 55.5 15.3 22.7 13.6
+7 63.1 54.6 14.8 22.5 13.2
+SST2 3 84.5 80.2 14.3 29.4 19.6
+5 88.6 82.1 18.5 26.8 20.4
+7 92.7 86.1 19.3 32.1 21.7
+Emo 3 58 35.2 11.5 14.6 7.9
+5 73.1 43.6 14.9 17.5 12.7
+7 79.2 47.4 11.7 15.8 11.3
+Poem 3 61 48.6 16.9 20.4 15.6
+5 62.9 51.4 18.1 23.3 17.2
+7 66.7 55.2 19.5 22.6 18.3
+AG 3 66.9 52.8 12.9 17.5 9.5
+5 73.2 57 13.6 19.4 11.9
+7 78 60.1 13.2 18.9 11.6
 Table 11: Perplexity scores for poisoned texts across different models and datasets.
@@
 AG 3.12 4.08 4.06 5.04
-Table 9: Full results for different number of examples in the demonstration. Focus on model Llama2-7B.
-Dataset num_examples Clean Random label Synonym Character Adv suffix
-Cola 3 63.2 59.2 16.5 21.8 12.8
-5 63.8 55.5 15.3 22.7 13.6
-7 63.1 54.6 14.8 22.5 13.2
-SST2 3 84.5 80.2 14.3 29.4 19.6
-5 88.6 82.1 18.5 26.8 20.4
-7 92.7 86.1 19.3 32.1 21.7
-Emo 3 58 35.2 11.5 14.6 7.9
-5 73.1 43.6 14.9 17.5 12.7
-7 79.2 47.4 11.7 15.8 11.3
-Poem 3 61 48.6 16.9 20.4 15.6
-5 62.9 51.4 18.1 23.3 17.2
-7 66.7 55.2 19.5 22.6 18.3
-AG 3 66.9 52.8 12.9 17.5 9.5
-5 73.2 57 13.6 19.4 11.9
-7 78 60.1 13.2 18.9 11.6
-A.3 Poisoned Text Examples
+Calibration and selection methods.
+We conduct additional experiments with two representative calibration and selection methods, contextual calibration (Zhao et al., 2021) and similarity sampling (Margatina et al., 2023), on the SST2 dataset with Llama2-7B.
+This is because the victims can apply these pre-processing methods.
+According to the results in Table12, calibration and selection methods can improve the clean accuracy.
+However, these methods are still vulnerable to the proposed poisoning attack, with a test accuracy of less than 20%.
+This may be because the contextual calibration only focuses on the potentially imbalanced distribution of the output label but does not consider the hidden states during calibration.
+The similarity selection selects demonstrations based on embedding similarity (Sentence-BERT as in (Margatina et al., 2023)), which may also be affected by the poisoning.
+Table 12: Test calibration and selection methods on SST2 dataset and Llama2-7B model.
+Clean Synonym Attack
+Random selection (baseline in our paper) 88.6 18.5
+Contextual calibration 90.4 19.2
+Similarity selection 91.6 17.4
+Scalability to context length and demonstration size.
+We additionally summarize the average token length of each dataset (to represent ‘context length’) used in the experiments.
+We use the tokenizer of Llama2-7B as the example and present results in Table 13.
+According to the above results, there is no obvious relationship between the attacking performance and the context length.
+This indicates that our method is robust to the context length.
+Table 13: Average token length of each dataset on Llama2-7B model.
+Dataset Avg token length attack/clean acc
+SST2 14.8 18.5/88.6
+COLA 10.4 15.3/63.8
+Poem 10.5 18.1/62.9
+Emo 18.6 11.9/73.1
+AG’s news 62.4 13.6/73.2
+Computation cost.
+We provide the average running time per example (in seconds) when poisoning dataset SST2 on model Llama2-7B with all three methods under different numbers of perturbations (1,3,5 perturbations).
+As shown in Table 14, adversarial suffix and synonym replacement can be time-consuming, while character replacement is much more efficient.
+This can be due to the size of the search space, as there are more tokens and words than characters.
+However, since the poisoning process is offline, attackers can take enough time to generate poisoned examples.
+Table 14: Computation cost, SST2, Llama2-7B
+Num of perturbation 1 3 5
+Adversarial suffix 127 369 571
+Char replacement 43 118 197
+Synonym replacement 71 206 343
+A.4 Poisoned Text Examples
 In this section, we provide some poisoned examples in Table 15 for human evaluation.
@@
 sentence: t)(h)eC( )worthy sucS(c)essor to a bettn(e)r toG(m)orrow
+A.5 Paraphrased texts
+To clarify about paraphrasing, we prompt the model to paraphrase the whole sentences.
+As a result, it is not a token-level replacement, and it is also hard to quantify how many tokens are changed.
+We present some examples in Table 16.
+According to the above table, the paraphrased sentence is different from the original text as it does not only change words but also changes the sentence structure.
+Besides, as in the third example, when adding some irrelevant content to the input, paraphrasing can remove it.
+This indicates that the paraphrasing defense is strong compared to attacks.
+Table 16: Poisoned texts paraphrased by GPT-4
+Original text Attacked text Paraphrased text
+Temperatures Dip in Europe, sending Price of Oil Up 4.8 The price of crude oil surged 4.8 percent in New York yesterday, the biggest rise in more than five months, on concern that cold weather in Europe will draw heating oil from the United States.
+Temperatures Dip in Europe, sent Price of Oil Up 4.8 The price of coarse oil surging 4.8 percent in New York yesterday, the biggest rise in more than five months, on stress that cold weather in Europe will draw heating oil from the unified states.
+European temperatures have fallen, leading to a 4.8% surging in coarse oil prices in New York yesterday.
+This marks the largest increase in over five months, driven by stresses that colder conditions in Europe will increase demand for heating oil from the unified states.
+More bodies found in Nablus US forces discovered more bodies under the northern city of Mosul on Friday, apparent victims of an intimidation campaign by insurgents against Iraq’s fledgling security forces.
+Additional bodies found in Nablus US forces uncovered more bodies under the northern city of Mosul on Friday, evident victims of an bullying campaign by insurgents against Iraq;s fledgling security forces.
+More bodies were discovered in Nablus as US forces unearthed additional victims in the northern city of Mosul on Friday, evidently targeted by insurgents in a campaign of bullying against Iraq’s emerging security forces.
+ETA Bombs Put Radical Basque Party in Tight Spot The armed Basque separatist group ETA showed it was not a spent force with Friday #39;s coordinated attacks at five Madrid petrol stations, but it also further isolated the guerrillas #39; closest political allies, Batasuna.
+ETA Bombs Put Radical Basque Party in Tight Spot The armed Basque separatist group ETA showed it was not a spent force with Friday #39;s coordinated attacks at five Madrid petrol stations, but it also further isolated the guerrillas #39; closest political allies, Batasuna.
+Gould}}\$,attempt The ETA’s bombings have placed its key political ally, the radical Basque party Batasuna, in a difficult position.
+By carrying out synchronized attacks at five Madrid petrol stations on Friday, the armed Basque separatist group demonstrated that it remains a potent force but also increasingly alienated Batasuna from others.